### v1

# Capillary $L_{p}$-Christoffel-Minkowski problem

Abstract.
We solve the capillary $L_{p}$-Christoffel–Minkowski problem in the half-space for $1<p<k+1$ in the class of even hypersurfaces. A crucial ingredient is a non-collapsing estimate that yields lower bounds for both the height and the capillary support function. Our result extends the capillary Christoffel–Minkowski existence result of [HIS25].
1. Introduction
The problem of prescribing area measures of convex hypersurfaces originates in the classical works of Christoffel [Chr65], Minkowski [Min97, Min03], Aleksandrov [Ale56], Nirenberg [Nir57] and Pogorelov [Pog52, Pog71], which established the modern interplay between convex geometry and fully nonlinear elliptic equations. In the smooth setting, the Christoffel–Minkowski problem seeks a smooth, strictly convex hypersurface whose $k$-th elementary symmetric function of the principal radii of curvature agrees with a given function on the sphere. This direction was further developed in the works of Firey and Berg [Fir67, Fir70, Ber69].
Over the past decades, the Christoffel–Minkowski problem has seen substantial progress. For the top-order case $k=n$, corresponding to the classical Minkowski problem, the situation is by now well understood: the seminal works of Cheng–Yau [CY76] and Caffarelli [Caf90a, Caf90b] provide an existence and regularity theory for the underlying fully nonlinear equation. For intermediate orders $1<k<n$, the picture is less complete, although [GM03, STW04] provide a far-reaching existence result for the Christoffel–Minkowski problem in the smooth setting. See also [BHO25, MU25] for the recent break-through in the rotationally symmetric case.
The $L_{p}$-extension of the Christoffel–Minkowski problem, introduced by Lutwak [Lut93] in the framework of the Brunn–Minkowski–Firey theory, replaces the classical area measures by their $L_{p}$ analogues and leads to the curvature equation
(1.1) $\sigma_{k}(\tau^{\sharp}[h])=h^{p-1}\phi\quad\text{on }\mathbb{S}^{n},$
for the support function $h$ of a smooth, strictly convex body/hypersurface, where $\tau^{\sharp}[h]=g^{-1}\cdot(\nabla^{2}h+hg)$ and $g$ denotes the standard metric on $\mathbb{S}^{n}$. The $L_{p}$-Minkowski problem has since been the subject of intensive study and has developed into a mature theory over a broad range of $p$; see [LO95, CW06, BLYZ13, HLYZ16, BBCY19, HXZ21, GLW22, LXYZ24] and also [CW00, BIS19, LWW20, CL21, BIS21b, BG23]. In contrast, for $k<n$ the situation is more fragmentary as the intermediate $L_{p}$-area measures remain, in general, much less understood. In the smooth case, however, and in particular for $p>1$ with even data on $\mathbb{S}^{n}$, one now has a well-developed set of results: existence, uniqueness and regularity of solutions, together with constant rank theorems ensuring strict convexity; see, for instance, [GM03, HMS04, GLM06, GMZ06, GX18, Iva19, BIS23a, BIS23b, HI24, Zha24, CH25] and [BIS21a, HLX24, LW24].
A natural question is how this picture changes in the presence of a boundary. In the capillary setting, one considers hypersurfaces in the half-space that meet a fixed supporting hyperplane at a prescribed contact angle $\theta\in(0,\pi/2)$. For the top-order case $k=n$, capillary versions of the $L_{p}$-Minkowski problem have been developed in a series of recent works. For $p\geq 1$, Mei, Wang and Weng solved the capillary $L_{p}$-Minkowski problem via the continuity method in [MWW25a, MWW25b]. For (n+1)<p<1-(n+1)<p<1, even solutions were constructed in [HI25] by means of an iterative scheme based on the curvature image operator, and a unified curvature flow approach was later introduced in [HHI25], treating the even capillary $L_{p}$-Minkowski problem for all p>(n+1)p>-(n+1).
For $k<n$, the capillary analogue of the Christoffel–Minkowski problem prescribes $\sigma_{k}(\tau^{\sharp}[s])$ on the capillary spherical cap $\mathcal{C}_{\theta}$ and couples the interior equation with a Robin boundary condition encoding the contact angle. This capillary analogue was solved in [HIS25], where the existence of smooth, strictly convex, $\theta$-capillary hypersurfaces was established under conditions on the prescribed function that are tailored to the applicability of a constant rank theorem.
The aim of this paper is to extend the work [HIS25] to the $L_{p}$-framework in the range $1<p<k+1$. In analogy with the closed case [GM03, GX18], we study the prescribed curvature equation
(1.2) $\sigma_{k}(\tau^{\sharp}[s])=s^{p-1}\phi\quad\text{in }\mathcal{C}_{\theta}$
for an even, positive, smooth function $\phi$ on $\mathcal{C}_{\theta}$, together with the capillary boundary condition
(1.3) $\nabla_{\mu}s=\cot\theta\,s\quad\text{on }\partial\mathcal{C}_{\theta}.$
Theorem 1.1.
Let $1<p<k+1$, $\theta\in(0,\pi/2)$, and $\phi\in C^{\infty}(\mathcal{C}_{\theta})$ be a positive function satisfying
(1.4) $\phi(-\zeta_{1},\ldots,-\zeta_{n},\zeta_{n+1})=\phi(\zeta_{1},\ldots,\zeta_{n},\zeta_{n+1})\quad\forall\zeta\in\mathcal{C}_{\theta},$
(1.5) $\nabla^{2}\phi^{-\frac{1}{p+k-1}}+g\phi^{-\frac{1}{p+k-1}}\geq 0\quad\text{in }\mathcal{C}_{\theta}$
and the boundary condition
(1.6) $\nabla_{\mu}\phi^{-\frac{1}{p+k-1}}\leq\cot\theta\,\phi^{-\frac{1}{p+k-1}}\quad\text{on }\partial\mathcal{C}_{\theta}.$
Then there exists a unique even, strictly convex, capillary hypersurface $\Sigma\subset\overline{\mathbb{R}^{n+1}_{+}}$ with contact angle $\theta$ whose capillary support function $s$ solves
(1.7) $\left\{\begin{aligned} \sigma_{k}(\tau^{\sharp}[s])&=s^{p-1}\phi&&\text{in }\mathcal{C}_{\theta},\\ \nabla_{\mu}s&=\cot\theta\,s&&\text{on }\partial\mathcal{C}_{\theta}.\end{aligned}\right$.
The paper is organized as follows. In Section 2 we recall the basic capillary geometry in the half-space and fix notation. Section 3 is devoted to non-collapsing estimates; i.e. a lower bound for the height of the hypersurface, both in the rotationally symmetric and in the general even case. In Section 4 we derive curvature and regularity estimates for solutions of (1.7). In Section 5 we prove a capillary constant rank theorem for our equation. Finally, in Section 6 we complete the proof of 1.1 by establishing existence and uniqueness.
2. Preliminaries
Let $\{e_{i}\}_{i=1}^{n+1}$ be the standard orthonormal basis of $\mathbb{R}^{n+1}$. Let
(2.1) $\mathbb{R}^{n+1}_{+}=\{x\in\mathbb{R}^{n+1}:x_{n+1}>0\}$
be the upper half-space with boundary $\partial\mathbb{R}^{n+1}_{+}=\{x_{n+1}=0\}$. The unit ball of $\mathbb{R}^{n+1}$ is denoted by $\mathbb{B}$, and we write $\mathbb{S}^{n}$ for the unit ball.
  1. (1)
    Support functions of convex bodies. For a bounded convex set $K\subset\mathbb{R}^{n+1}$, the support function $h_{K}:\mathbb{S}^{n}\to\mathbb{R}$ is defined as
    (2.2) $h_{K}(u)=\sup\{\langle x,u\rangle:x\in K\},\quad u\in\mathbb{S}^{n}.$
    When no confusion can arise, we simply write $h:=h_{K}$.
  2. (2)
    Area measures in $\mathbb{R}^{n+1}$. Let $K\subset\mathbb{R}^{n+1}$ a bounded convex set and $k\in\{0,\dots,n\}$. The $k$-th area measure $S_{k}(K,\cdot)$ is the finite Borel measure on $\mathbb{S}^{n}$ appearing in the classical local Steiner formula. If $K$ is smooth and strictly convex with principal radii of curvature $\lambda_{1},\dots,\lambda_{n}$ at a point with outer unit normal $u\in\mathbb{S}^{n}$, then
    (2.3) $dS_{k}(K,u)=\frac{1}{\binom{n}{k}}\sigma_{k}(\lambda_{1},\dots,\lambda_{n})\,d\sigma(u),$
    where $d\sigma$ denotes the spherical Lebesgue measure on $\mathbb{S}^{n}$. If $L\subset\mathbb{R}^{n+1}$ is an $m$-dimensional linear subspace and $K\subset L$, we write $S_{k}^{L}(K,\cdot)$ for the $k$-th area measure of $K$ viewed as a convex set in $L$.
  3. (3)
    Hausdorff measure and subspheres. For any integer $d\geq 1$, we write $\mathcal{H}^{d}$ for the $d$-dimensional Hausdorff measure, and
    (2.4) $\mathbb{S}^{d}=\{x\in\mathbb{R}^{d+1}:|x|=1\}$
    for the unit sphere in $\mathbb{R}^{d+1}$. For a linear subspace $L\subset\mathbb{R}^{n+1}$ of dimension $m$, we identify $\mathbb{S}^{n}\cap L$ with the unit sphere in $L$. We also write
    (2.5) $\mathbb{S}^{n}_{\theta}=\{x\in\mathbb{S}^{n}:\langle x,e_{n+1}\rangle\geq\cos\theta\},\quad\mathcal{C}_{\theta}=\mathbb{S}^{n}_{\theta}-\cos\theta\,e_{n+1}.$
    Integrals of the form
    (2.6) $\int_{\mathbb{S}^{n}}f,\quad\int_{\mathbb{S}^{n}_{\theta}}f,\quad\int_{\mathcal{C}_{\theta}}f,\quad\int_{\mathbb{S}^{n}\cap L}f,\quad\int_{\mathbb{S}^{d}}f,$
    are always understood with respect to the restriction of the appropriate Hausdorff measure (thus, $\mathcal{H}^{n}$ on $\mathbb{S}^{n}$, $\mathbb{S}^{n}_{\theta}$ and $\mathcal{C}_{\theta}$, $\mathcal{H}^{m-1}$ on $\mathbb{S}^{n}\cap L$, and $\mathcal{H}^{d}$ on $\mathbb{S}^{d}$). We also write
    (2.7) $\omega_{d}=\mathcal{H}^{d}(\mathbb{S}^{d}),$
    so that $\omega_{d}$ is the surface area of $\mathbb{S}^{d}$.
Definition 2.1.
A smooth, compact, connected, orientable hypersurface $\Sigma\subset\overline{\mathbb{R}^{n+1}_{+}}$ with $\mathrm{int}(\Sigma)\subset\mathbb{R}^{n+1}_{+}$ and $\partial\Sigma\subset\partial\mathbb{R}^{n+1}_{+}$ is called a capillary hypersurface with contact angle $\theta\in(0,\pi)$ if
(2.8) $\langle\nu,e_{n+1}\rangle=\cos\theta\quad\text{on }\partial\Sigma,$
where $\nu$ is the outer unit normal of $\Sigma$.
The model capillary surface is
(2.9) $\mathcal{C}_{\theta}=\{\zeta\in\overline{\mathbb{R}^{n+1}_{+}}:|\zeta+\cos\theta\,e_{n+1}|=1\}.$
Via the translation
(2.10) $T(\zeta)=\zeta+\cos\theta\,e_{n+1},$
we may identify $\mathcal{C}_{\theta}$ with $\mathbb{S}^{n}_{\theta}$.
We also define
(2.11) $\mathcal{C}_{\theta,r}=\left\{\zeta\in\overline{\mathbb{R}_{+}^{n+1}}\mid|\zeta+r\cos\theta\,e_{n+1}|=r\right\}.$
Note that the radius of $\partial\mathcal{C}_{\theta,r}$ is $r\sin\theta$.
We call $\Sigma$ strictly convex if the enclosed region $\widehat{\Sigma}$ is a convex body (i.e. compact, convex, with non-empty interior) and the second fundamental form of $\Sigma$ is positive definite. For a strictly convex capillary hypersurface $\Sigma$, the capillary Gauss map is defined as
(2.12) $\tilde{\nu}=\nu-\cos\theta\,e_{n+1}:\Sigma\to\mathcal{C}_{\theta}.$
This is a diffeomorphism onto the capillary spherical cap, see [MWWX25, Lem. 2.2].
Definition 2.2.
Let $\Sigma$ be a strictly convex, capillary hypersurface. The capillary support function $s:\mathcal{C}_{\theta}\to\mathbb{R}$ of $\Sigma$ is defined by
(2.13) $s(\zeta)=\langle\tilde{\nu}^{-1}(\zeta),\zeta+\cos\theta\,e_{n+1}\rangle.$
For the model cap $\mathcal{C}_{\theta}$, the capillary support function is
(2.14) $\ell(\zeta)=\sin^{2}\theta-\cos\theta\langle\zeta,e_{n+1}\rangle.$
On $\mathcal{C}_{\theta}$ we also write $g$ for the round metric, $\nabla$ for its Levi-Civita connection and $\nabla^{2}$ for the covariant Hessian. For a function $f\in C^{2}(\mathcal{C}_{\theta})$ we set
(2.15) $\tau[f]=\nabla^{2}f+f\,g,\quad\tau^{\sharp}[f]=g^{-1}\cdot\tau[f],$
so that $\tau^{\sharp}[f]$ is a symmetric endomorphism of $T\mathcal{C}_{\theta}$. Its eigenvalues are denoted by $\lambda_{1},\dots,\lambda_{n}$, and $\sigma_{k}(\tau^{\sharp}[f])$ means $\sigma_{k}(\lambda_{1},\dots,\lambda_{n})$. We also write $\nabla_{\mu}f$ for the covariant derivative in the direction of the outward unit conormal $\mu$ along $\partial\mathcal{C}_{\theta}$.
For a symmetric matrix $A=(a_{ij})$ with eigenvalues $\lambda_{1},\dots,\lambda_{n}$ we write
(2.16) $\sigma_{k}^{ij}(A)=\frac{\partial\sigma_{k}}{\partial a_{ij}}(A),$
and for $F=\sigma_{k}^{1/k}$ we set
(2.17) $F^{ij}(A)=\frac{\partial F}{\partial a_{ij}}(A)=\frac{1}{k}\,\sigma_{k}(A)^{\frac{1}{k}-1}\sigma_{k}^{ij}(A).$
When $A=\tau^{\sharp}[s]$ for some function $s$ on $\mathcal{C}_{\theta}$, we abbreviate $\sigma_{k}^{ij}(\tau^{\sharp}[s])$ and $F^{ij}(\tau^{\sharp}[s])$ by $\sigma_{k}^{ij}$ and $F^{ij}$, respectively.
Writing points of $\mathbb{R}^{n+1}$ as $x=(x_{1},\dots,x_{n},x_{n+1})$, let $\mathcal{R}$ denote the reflection
(2.18) $\mathcal{R}(x_{1},\dots,x_{n},x_{n+1})=(-x_{1},\dots,-x_{n},x_{n+1}).$
A function $\varphi:\mathcal{C}_{\theta}\to\mathbb{R}$ is called even if $\varphi\circ\mathcal{R}=\varphi$, and we say that a capillary hypersurface $\Sigma$ (or its capillary support function $s$) is even if
(2.19) $x\in\Sigma\implies\mathcal{R}(x)\in\Sigma.$
Definition 2.3.
Let $k\in\{0,\dots,n\}$. Let $s_{0},\dots,s_{n}\in C^{\infty}(\mathcal{C}_{\theta})$ be capillary support functions. Denote by $Q_{k}$ the linear polarization of $\sigma_{k}$ on symmetric endomorphisms of $T\mathcal{C}_{\theta}$, i.e. the unique symmetric multilinear map such that
(2.20) $Q_{k}(A,\dots,A)=\frac{\sigma_{k}(A)}{\binom{n}{k}}\quad\text{for every symmetric endomorphism }A.$
The capillary mixed volume of $s_{0},\dots,s_{k}$ is defined by
(2.21) $V(s_{0},\dots,s_{k},\underbrace{\ell,\ldots,\ell}_{n-k\ \text{times}})=\frac{1}{n+1}\int_{\mathcal{C}_{\theta}}s_{0}Q_{k}\bigl(\tau^{\sharp}[s_{1}],\dots,\tau^{\sharp}[s_{k}],\underbrace{\tau^{\sharp}[\ell],\dots,\tau^{\sharp}[\ell]}_{n-k\ \text{times}}\bigr).$
In particular, one has
(2.22) $V(s_{0},\underbrace{s,\dots,s}_{k\ \text{times}},\ell,\dots,\ell)=\frac{1}{n+1}\int_{\mathcal{C}_{\theta}}s_{0}\frac{\sigma_{k}\bigl(\tau^{\sharp}[s]\bigr)}{\binom{n}{k}}$.
$x_{n+1}=0$$x_{n+1}=-t\cos\theta$$\Sigma_{t}$$\Sigma$$\Sigma-t\cos\theta\,e_{n+1}$
Figure 1. Capillary vs. classical outer parallel hypersurfaces
Theorem 2.4.
Let $\Sigma\subset\overline{\mathbb{R}^{n+1}_{+}}$ be a strictly convex, $\theta$-capillary hypersurface. For $t>0$ define
(2.23) $\phi_{t}:\Sigma\to\overline{\mathbb{R}^{n+1}_{+}},\quad\phi_{t}(x)=x+t\tilde{\nu}(x).$
Then $\Sigma_{t}:=\phi_{t}(\Sigma)$ is a strictly convex, $\theta$-capillary hypersurface. Moreover, the (standard) outer parallel convex body
(2.24) $K_{t}=\bigl(\widehat{\Sigma}-t\cos\theta\,e_{n+1}\bigr)+t\,\mathbb{B},$
and the capillary outer parallel convex body are related via $\widehat{\Sigma_{t}}=K_{t}\cap\overline{\mathbb{R}^{n+1}_{+}}$. In addition, we have $\Sigma_{t}=\Sigma+t\mathcal{C}_{\theta}$.
Proof.
Let $P=\{x_{n+1}=0\}$. Define $f(x)=\langle x,e_{n+1}\rangle=x_{n+1}$ and
(2.25) $g(x)=\langle\tilde{\nu}(x),e_{n+1}\rangle=\langle\nu(x),e_{n+1}\rangle-\cos\theta.$
Then
(2.26) $\langle\phi_{t}(x),e_{n+1}\rangle=f(x)+t\,g(x).$
Step 1. On $\partial\Sigma$ we have $f=0$ (since $\partial\Sigma\subset P$) and $g=0$ (by capillarity), hence $(f+tg)(x)=0$ for $x\in\partial\Sigma$, i.e. $\phi_{t}(\partial\Sigma)\subset P$. Moreover, since $\nu(\operatorname{int}(\Sigma))\subset\operatorname{int}(\mathbb{S}^{n}_{\theta})$, $f+tg>0$ on $\operatorname{int}(\Sigma)$ for any $t>0$.
Step 2. Let $x\in\Sigma$ and choose an orthonormal basis $\{e_{1},\dots,e_{n}\}$ of $T_{x}\Sigma$ consisting of principal directions, so that
(2.27) ${}^{\Sigma}\nabla_{e_{i}}\nu=\kappa_{i}e_{i},\quad i=1,\dots,n,$
with principal curvatures $\kappa_{i}$. We have $^{\Sigma}\nabla_{e_{i}}\tilde{\nu}={}^{\Sigma}\nabla_{e_{i}}\nu$, and therefore
(2.28) $d\phi_{t}(e_{i})=e_{i}+t\,{}^{\Sigma}\nabla_{e_{i}}\tilde{\nu}=e_{i}+t\,{}^{\Sigma}\nabla_{e_{i}}\nu=(1+t\kappa_{i})e_{i}.$
Thus $d\phi_{t}(T_{x}\Sigma)$ is spanned by $\{e_{1},\dots,e_{n}\}$, $\phi_{t}$ is a smooth immersion, and the oriented unit normal of $\Sigma_{t}=\phi_{t}(\Sigma)$ at $y=\phi_{t}(x)$ equals $\nu(x)$, i.e.
(2.29) $\nu_{t}(y)=\nu(x)\quad\text{for }y=\phi_{t}(x).$
Next we show that $\phi_{t}$ is an injective immersion and thus an embedding. Assume $\phi_{t}(x)=\phi_{t}(x^{\prime})$ for some $x,x^{\prime}\in\Sigma$. Then
(2.30) $x+t(\nu(x)-\cos\theta\,e_{n+1})=x^{\prime}+t(\nu(x^{\prime})-\cos\theta\,e_{n+1}),$
hence
(2.31) $x-x^{\prime}=t\bigl(\nu(x^{\prime})-\nu(x)\bigr).$
Taking the inner product with $\nu(x)$ gives
(2.32) $\langle x-x^{\prime},\nu(x)\rangle=t\bigl(\langle\nu(x^{\prime}),\nu(x)\rangle-1\bigr).$
Since $\widehat{\Sigma}$ is convex and $\nu(x)$ is the outer normal vector to $\Sigma$ at $x$,
(2.33) $\langle x^{\prime}-x,\nu(x)\rangle\leq 0.$
On the other hand, $\langle\nu(x^{\prime}),\nu(x)\rangle\leq 1$, hence by (2.32)
(2.34) $\langle x-x^{\prime},\nu(x)\rangle=0\quad\text{and}\quad\langle\nu(x^{\prime}),\nu(x)\rangle=1.$
Therefore $\nu(x^{\prime})=\nu(x)$. Since $\Sigma$ is strictly convex, the Gauss map $\nu:\Sigma\to\mathbb{S}_{\theta}^{n}$ is injective, hence $x^{\prime}=x$.
Step 3. If $y=\phi_{t}(x)$ with $x\in\partial\Sigma$, then by step 1 and step 2 we have $y\in P$ and $\nu_{t}(y)=\nu(x)$. Therefore,
(2.35) $\langle\nu_{t}(y),e_{n+1}\rangle=\langle\nu(x),e_{n+1}\rangle=\cos\theta,$
so $\Sigma_{t}$ meets $P$ with the same contact angle $\theta$.
Step 4. Let $x\in\Sigma$ and set
(2.36) $y=\phi_{t}(x)=x+t(\nu(x)-\cos\theta\,e_{n+1}).$
We claim that $y\in\partial K_{t}$ and that $\nu(x)$ is an outer normal of $K_{t}$ at $y$.
Note that $y\in K_{t}$. Since $\nu(x)$ is an outer unit normal of the convex body $\widehat{\Sigma}$ at $x$, we have
(2.37) $\langle z-x,\nu(x)\rangle\leq 0\quad\forall\,z\in\widehat{\Sigma}$.
Let $w\in K_{t}$. Then for some $z\in\widehat{\Sigma}$ and $b\in\mathbb{B}$:
(2.38) $w=(z-t\cos\theta\,e_{n+1})+tb.$
Moreover, we have
(2.39) $\langle w-y,\nu(x)\rangle=\langle z-x,\nu(x)\rangle+t\langle b,\nu(x)\rangle-t.$
Using (2.37), we obtain
(2.40) $\langle w-y,\nu(x)\rangle\leq 0\quad\forall\,w\in K_{t}.$
Hence we must have $y\in\partial K_{t}$ and $\nu(x)$ is an outer normal of $K_{t}$ at $y$.
Step 5. Let $L_{t}=K_{t}\cap\overline{\mathbb{R}^{n+1}_{+}}$. We prove
(2.41) $\Sigma_{t}=\phi_{t}(\Sigma)\subset\partial L_{t}.$
If $x\in\operatorname{int}(\Sigma)$, then by step 1, we have
(2.42) $y=\phi_{t}(x)\in\mathbb{R}^{n+1}_{+}.$
Together with $y\in\partial K_{t}$ (by step 4) this implies $y\in\partial L_{t}\setminus P$.
If $x\in\partial\Sigma$, then by step 1, $\phi_{t}(\partial\Sigma)\subset P$, so $y\in P$. Let $x_{j}\in\operatorname{int}(\Sigma)$ be any sequence with $x_{j}\to x$. Set $y_{j}:=\phi_{t}(x_{j})$. By continuity, $y_{j}\to y$. Since $y_{j}\in\partial L_{t}$ the limit point $y$ belongs to $\partial L_{t}\cap P$.
Step 6. We prove $\overline{\partial L_{t}\setminus P}=\Sigma_{t}$. By steps 1, 2 and 5,
(2.43) $\operatorname{int}(\Sigma_{t})=\phi_{t}(\operatorname{int}(\Sigma))\subset\partial L_{t}\setminus P\implies\Sigma_{t}\subset\overline{\partial L_{t}\setminus P}.$
It remains to prove $\partial L_{t}\setminus P\subset\Sigma_{t}\setminus P$.
Let $y\in\partial L_{t}\setminus P$. Then $y\in\partial K_{t}$. Suppose $u\in\mathbb{S}^{n}$ is an outer unit normal to $K_{t}$ at $y$, i.e.
(2.44) $\langle w-y,u\rangle\leq 0\quad\forall\,w\in K_{t}.$
We may write
(2.45) $y=(x-t\cos\theta\,e_{n+1})+tb,\quad x\in\widehat{\Sigma},\quad b\in\mathbb{B}.$
We claim that $b=u$, $x\in\partial\widehat{\Sigma}$, and $u$ is an outer normal of $\widehat{\Sigma}$ at $x$.
Indeed, take any $x_{0}\in\widehat{\Sigma}$ and any $c\in\mathbb{B}$, and set
(2.46) $w=(x_{0}-t\cos\theta\,e_{n+1})+tc\in K_{t}.$
Plugging this $w$ and (2.45) into (2.44) gives
(2.47) $0\geq\langle w-y,u\rangle=\langle x_{0}-x,u\rangle+t\langle c-b,u\rangle.$
With $x_{0}=x$ and $c=u$,
(2.48) $1\leq\langle b,u\rangle\implies b=u.$
Now with $c=b=u$, the inequality becomes
(2.49) $0\geq\langle x_{0}-x,u\rangle\quad\forall\,x_{0}\in\widehat{\Sigma}$,
so $x\in\partial\widehat{\Sigma}$ and $u$ is an outer normal of $\widehat{\Sigma}$ at $x$.
Since $y_{n+1}>0$ and $t>0$, we have from (2.45) (with $b=u$)
(2.50) $y_{n+1}=x_{n+1}-t\cos\theta+tu_{n+1}>0.$
This implies that $x_{n+1}>0$, $x\in\Sigma$ and $u=\nu(x)$ (otherwise, if $x_{n+1}=0$, then we would have $u_{n+1}\leq\cos\theta$ and hence $y_{n+1}\leq 0$). Substituting $b=u=\nu(x)$ into (2.45) yields
(2.51) $y=x+t(\nu(x)-\cos\theta\,e_{n+1})=\phi_{t}(x)\in\Sigma_{t}\setminus P.$
Step 7. By the previous steps, $\Sigma_{t}$ is a strictly convex (i.e. the enclosed region $\widehat{\Sigma_{t}}$ is a convex body and the second fundamental form of $\Sigma_{t}$ is positive definite), $\theta$-capillary hypersurface, and for each point $x\in\Sigma$, the outward unit normal at the point $\phi_{t}(x)\in\Sigma_{t}$ is $\nu(x)$. Let $\zeta=\nu(x)-\cos\theta\,e_{n+1}$. Then
(2.52) $s_{\Sigma_{t}}(\zeta)$ $=\langle\phi_{t}(x),\nu(x)\rangle$
$=\langle x+t(\nu(x)-\cos\theta e_{n+1}),\nu(x)\rangle$
$=s_{\Sigma}(\zeta)+t\ell(\zeta),$
Since $\Sigma_{t}$ has the same capillary support function as $\Sigma+t\mathcal{C}_{\theta}$, we conclude
(2.53) $\Sigma_{t}=\Sigma+t\mathcal{C}_{\theta}.$
For $\rho>0$ and a Borel set $\omega\subset\mathcal{C}_{\theta}$ we define the local outer parallel set of $\widehat{\Sigma}$ in the directions of $\omega$ by
(2.54) $B_{\rho,\theta}(\widehat{\Sigma},\omega)$ $=\left\{y\in\overline{\mathbb{R}^{n+1}_{+}}:\begin{array}[]{l}\exists x\in\Sigma,\,0<t<\rho,\,\text{s.t.}\\ y=x+t\tilde{\nu}(x),\,\tilde{\nu}(x)\in\omega\end{array}\right\}.$
Lemma 2.5.
Let $\Sigma\subset\overline{\mathbb{R}^{n+1}_{+}}$ be a strictly convex $\theta$-capillary hypersurface with principal curvatures $\kappa=(\kappa_{1},\dots,\kappa_{n})$ and area element $d\mu$. Then, for every Borel set $\omega\subset\mathcal{C}_{\theta}$ and every $\rho>0$,
(2.55) $\operatorname{vol}\bigl(B_{\rho,\theta}(\widehat{\Sigma},\omega)\bigr)=\sum_{j=0}^{n}\frac{\rho^{n+1-j}}{n+1-j}\int_{\Sigma\cap\tilde{\nu}^{-1}(\omega)}\bigl(1-\cos\theta\langle\nu,e_{n+1}\rangle\bigr)\sigma_{n-j}(\kappa)\,d\mu.$
Proof.
Let
(2.56) $\Phi:\Sigma\times(0,\infty)\to\overline{\mathbb{R}^{n+1}_{+}},\quad\Phi(x,t)=x+t\tilde{\nu}(x).$
By 2.4, $\Phi$ maps $\Sigma$ to strictly convex, $\theta$-capillary hypersurfaces.
For a given Borel set $\omega\subset\mathcal{C}_{\theta}$, the definition (2.54) gives
(2.57) $B_{\rho,\theta}(\widehat{\Sigma},\omega)=\Phi\Bigl(\bigl(\Sigma\cap\tilde{\nu}^{-1}(\omega)\bigr)\times(0,\rho)\Bigr).$
Hence
(2.58) $\operatorname{vol}\bigl(B_{\rho,\theta}(\widehat{\Sigma},\omega)\bigr)=\int_{\Sigma\cap\tilde{\nu}^{-1}(\omega)}\int_{0}^{\rho}J(x,t)\,dt\,d\mu(x),$
where $J(x,t)$ denotes the Jacobian of $\Phi$ at $(x,t)$.
Set $e=-e_{n+1}$. We have
(2.59) $J(x,t)$ $=\langle\tilde{\nu}(x),\nu(x)\rangle\prod_{i=1}^{n}(1+t\kappa_{i}(x))$
$=\bigl(1+\cos\theta\,\langle\nu(x),e\rangle\bigr)\prod_{i=1}^{n}(1+t\kappa_{i}(x))$
$=\bigl(1+\cos\theta\,\langle\nu(x),e\rangle\bigr)\sum_{j=0}^{n}\sigma_{n-j}(\kappa(x))t^{n-j}.$
Inserting this into (2.58) and integrating in $t$ yields
(2.60) $\operatorname{vol}\bigl(B_{\rho,\theta}(\widehat{\Sigma},\omega)\bigr)$ $=\int_{\Sigma\cap\tilde{\nu}^{-1}(\omega)}\int_{0}^{\rho}\bigl(1+\cos\theta\langle\nu,e\rangle\bigr)\sum_{j=0}^{n}\sigma_{n-j}(\kappa)t^{n-j}\,dt\,d\mu$
$=\sum_{j=0}^{n}\frac{\rho^{n+1-j}}{n+1-j}\int_{\Sigma\cap\tilde{\nu}^{-1}(\omega)}\bigl(1+\cos\theta\langle\nu,e\rangle\bigr)\sigma_{n-j}(\kappa)\,d\mu.$
Definition 2.6.
Let $\theta\in(0,\pi/2)$ and suppose $\Sigma\subset\mathbb{R}^{n+1}_{+}$ is a strictly convex, $\theta$-capillary hypersurface. For a Borel set $\omega\subset\mathcal{C}_{\theta}$ we define the capillary $k$-th area measure of $\widehat{\Sigma}$ over $\omega$ by
(2.61) $S_{k,\theta}(\widehat{\Sigma},\omega)=\binom{n}{k}^{-1}\int_{\tilde{\nu}^{-1}(\omega)}(1-\cos\theta\,\langle\nu,e_{n+1}\rangle)\sigma_{n-k}(\kappa)\,d\mu.$
[…]
(2.62) $dS_{k,\theta}(\widehat{\Sigma},\xi)=\binom{n}{k}^{-1}\ell(\xi)\sigma_{k}\bigl(\tau^{\sharp}[s](\xi)\bigr)\,d\mathcal{H}^{n}(\xi),\quad\xi\in\mathcal{C}_{\theta}.$
In particular,
(2.63) $S_{k,\theta}(\widehat{\Sigma},\omega)=\binom{n}{k}^{-1}\int_{\omega}\ell(\xi)\sigma_{k}\bigl(\tau^{\sharp}[s](\xi)\bigr)\,d\mathcal{H}^{n}(\xi),\quad\omega\subset\mathcal{C}_{\theta}\ \text{Borel}.$
Remark 2.7.
[…]
(2.64) $S_{k,\theta}(\widehat{\Sigma},\omega)=\ell\,S_{k}(\widehat{\Sigma},T\omega),$
so on such sets the capillary $k$-th area measure agrees (up to the weight $\ell$) with the restriction of the classical $k$-th area measure of $\widehat{\Sigma}$.
[…]
Theorem 2.8.
Assume $\Sigma$ is the graph of a convex function $f\in C^{1}(\Omega)$ on a bounded, closed convex set $\Omega$ with $f=0$ on $\partial\Omega$. Then for all $x^{\prime}\in\Omega$,
(2.65) $|Df(x^{\prime})|\leq\tan\theta,\quad|f(x^{\prime})|\leq\tan\theta\,\operatorname{dist}(x^{\prime},\partial\Omega).$
Set $H=\|f\|_{C(\Omega)}=\max\limits_{x\in\Sigma}\langle x,e_{n+1}\rangle$. If $\Sigma$ is even, then
(2.66) $B_{\frac{H}{\tan\theta}}(0)\subset\Omega,\quad\widehat{\mathcal{C}}_{\theta,\frac{H}{\tan\theta\sin\theta}}\subset\widehat{\Sigma}$.
Proof.
Since $\Sigma$ is the graph of $f$, we can write
(2.67) $-\Sigma=\{(x^{\prime},f(x^{\prime})):x^{\prime}\in\Omega\},\quad f\leq 0,\quad f=0\ \text{on }\partial\Omega.$
At a boundary point $x^{\prime}_{0}\in\partial\Omega$, the upward unit normal of the graph of $f$ is
(2.68) $\nu(x_{0}^{\prime})=\frac{1}{\sqrt{1+|Df(x^{\prime}_{0})|^{2}}}\bigl(-Df(x^{\prime}_{0}),1\bigr).$
By the capillary condition,
(2.69) $\langle\nu,e_{n+1}\rangle=\cos\theta\quad\Longrightarrow\quad\frac{1}{\sqrt{1+|Df|^{2}}}=\cos\theta,$
hence $|Df(x^{\prime}_{0})|=\tan\theta$ for every $x^{\prime}_{0}\in\partial\Omega$. Since $f$ is convex and $\Omega$ is bounded and convex, the maximum of $|Df|$ over $\Omega$ is attained on $\partial\Omega$, so
(2.70) $|Df|\leq\tan\theta\quad\text{in }\Omega.$
Let $x^{\prime}\in\Omega$ and choose $y^{\prime}\in\partial\Omega$ such that
(2.71) $|x^{\prime}-y^{\prime}|=\operatorname{dist}(x^{\prime},\partial\Omega).$
Set
(2.72) $\xi=\frac{x^{\prime}-y^{\prime}}{|x^{\prime}-y^{\prime}|},\quad g(t)=f\bigl(y^{\prime}+t\xi\bigr),\quad t\in[0,|x^{\prime}-y^{\prime}|].$
Then $g$ is convex, $g(0)=f(y^{\prime})=0$ and $g(|x^{\prime}-y^{\prime}|)=f(x^{\prime})\leq 0$. Using (2.70), we have $|g^{\prime}(t)|\leq\tan\theta$, hence
(2.73) $|f(x^{\prime})|\leq\int_{0}^{|x^{\prime}-y^{\prime}|}|g^{\prime}(t)|\,dt\leq\tan\theta\,\operatorname{dist}(x^{\prime},\partial\Omega).$
This gives the second inequality.
Assume now that $\Sigma$ is even. Then $f$ is even, i.e.
(2.74) $f(-x^{\prime})=f(x^{\prime})\quad\forall x^{\prime}\in\Omega,$
and $\Omega$ is origin-symmetric. For any $x^{\prime}\in\Omega$, convexity and evenness give
(2.75) $f(0)\leq\tfrac{1}{2}f(x^{\prime})+\tfrac{1}{2}f(-x^{\prime})=f(x^{\prime}),$
so $f(0)=\min_{\Omega}f=-H$.
Applying the distance estimate at $x^{\prime}=0$ yields
(2.76) $H=-f(0)\leq\operatorname{dist}(0,\partial\Omega)\tan\theta,$
and therefore
(2.77) $B_{\frac{H}{\tan\theta}}(0)\subset\Omega.$
To prove the last claim, consider the ($\theta$-capillary) cone in $\mathbb{R}^{n+1}$ with apex at (0,H)(0,-H) and base $B_{\frac{H}{\tan\theta}}(0)\subset\Omega$:
(2.78) $\mathcal{K}^{-}=\Bigl\{(x^{\prime},x_{n+1}):|x^{\prime}|\leq\frac{H}{\tan\theta},\ -H\leq x_{n+1}\leq-H+\tan\theta\,|x^{\prime}|\Bigr\}.$
The lateral boundary of $\mathcal{K}^{-}$ is the graph of
(2.79) $g(x^{\prime})=-H+\tan\theta\,|x^{\prime}|\quad\text{on }B_{\frac{H}{\tan\theta}}(0).$
Using (2.70) and f(0)=Hf(0)=-H, for $|x^{\prime}|\leq H/\tan\theta$ we have
(2.80) $f(x^{\prime})-f(0)=\int_{0}^{1}\langle Df(tx^{\prime}),x^{\prime}\rangle\,dt\leq\tan\theta\,|x^{\prime}|,$
hence
(2.81) $f(x^{\prime})\leq-H+\tan\theta\,|x^{\prime}|=g(x^{\prime}).$
Thus, for every such $x^{\prime}$,
(2.82) $\{x_{n+1}:g(x^{\prime})\leq x_{n+1}\leq 0\}\subset\{x_{n+1}:f(x^{\prime})\leq x_{n+1}\leq 0\}.$
Therefore, $\mathcal{K}^{-}\subset\widehat{\Sigma^{-}}$, where
(2.83) $\widehat{\Sigma^{-}}=\{(x^{\prime},x_{n+1}):x^{\prime}\in\Omega,\ f(x^{\prime})\leq x_{n+1}\leq 0\}$
is the region between the graph of $f$ and $\{x_{n+1}=0\}$.
Since the cap $\widehat{\mathcal{C}}_{\theta,\frac{H}{\tan\theta\sin\theta}}$ is contained in $\mathcal{K}^{-}$, we obtain
(2.84) $\widehat{\mathcal{C}}_{\theta,\frac{H}{\tan\theta\sin\theta}}\subset\widehat{\Sigma}$.
This completes the proof. ∎
3. Non-collapsing Estimates
Let $\theta\in(0,\pi/2)$, $p\in(1,k+1)$ and $q\in[1,p]$. Let $\Sigma$ be an even, strictly convex, $\theta$-capillary hypersurface whose capillary support function $s>0$ solves
(3.1) $s^{1-q}\sigma_{k}(\tau^{\sharp}[s])=\phi\quad\text{in }\mathcal{C}_{\theta},$
with the prescribed function $\phi\in C^{\infty}(\mathcal{C}_{\theta})$. Assume $\phi_{0}\leq\phi\leq\phi_{1}$ with the constants $0<\phi_{0}<1<\phi_{1}$.
Lemma 3.1.
Let $s$ satisfy (3.1). Then there exists a constant
(3.2) $C_{0}=C_{0}(n,k,p,\theta,\phi_{0},\phi_{1})>1$
such that
(3.3) $s\leq C_{0}\quad\text{on }\mathcal{C}_{\theta}.$
Proof.
Throughout the proof, constants depend only on $(n,k,p,\theta,\phi_{0},\phi_{1})$.
Integrating by parts (cf. [MWWX25, Cor. 2.10]) and using the Newton–Maclaurin inequality yields
(3.4) $c_{k}^{\prime}\int_{\mathcal{C}_{\theta}}s^{1+\frac{k-1}{k}(q-1)}\phi^{\frac{k-1}{k}}\leq\int_{\mathcal{C}_{\theta}}s\sigma_{k-1}=c_{k}\int_{\mathcal{C}_{\theta}}\ell\sigma_{k}=c_{k}\int_{\mathcal{C}_{\theta}}\ell\phi s^{q-1}.$
We can rewrite (3.4) as
(3.5) $\int_{\mathcal{C}_{\theta}}s^{\alpha(q)}\leq C_{1}\int_{\mathcal{C}_{\theta}}s^{\beta(q)},$
where
(3.6) $\beta(q)=q-1,\quad\alpha(q)=1+\frac{k-1}{k}(q-1)=1+\frac{k-1}{k}\beta(q),$
and $C_{1}=C_{1}(n,k,\theta,\phi_{0},\phi_{1})>1$. Since $q\in[1,p]$ with $1<p<k+1$, we have
(3.7) $0\leq\beta(q)\leq p-1<k,\quad 1\leq\alpha(q)\leq 1+\frac{k-1}{k}(p-1)<k.$
Assume $1<q\leq p$. Then $\beta(q)>0$ and $0<\beta(q)<\alpha(q)$, and by Hölder’s inequality,
(3.8) $\Bigl(\int_{\mathcal{C}_{\theta}}s^{\beta(q)}\Bigr)^{\frac{\alpha(q)}{\beta(q)}}\leq|\mathcal{C}_{\theta}|^{\frac{\alpha(q)}{\beta(q)}-1}\int_{\mathcal{C}_{\theta}}s^{\alpha(q)}.$
Combining with (3.5) we obtain
(3.9) $\int_{\mathcal{C}_{\theta}}s^{\beta(q)}\leq|\mathcal{C}_{\theta}|C_{1}^{\frac{\beta(q)}{\alpha(q)-\beta(q)}}.$
Note that
(3.10) $\alpha(q)-\beta(q)=1-\frac{1}{k}\beta(q),\quad\frac{\beta(q)}{\alpha(q)-\beta(q)}=\frac{\beta(q)}{1-\beta(q)/k}.$
Since $\beta(q)\in[0,p-1]$, we have
(3.11) $\frac{\beta(q)}{1-\beta(q)/k}\leq E_{p}=\frac{k(p-1)}{k+1-p}.$
Thus for $1<q\leq p$,
(3.12) $\int_{\mathcal{C}_{\theta}}s^{q-1}=\int_{\mathcal{C}_{\theta}}s^{\beta(q)}\leq|\mathcal{C}_{\theta}|C_{1}^{E_{p}}.$
Choosing the constant larger if necessary,
(3.13) $C_{2}=C_{2}(n,k,p,\theta,\phi_{0},\phi_{1})$
we have for all $q\in[1,p]$:
(3.14) $\int_{\mathcal{C}_{\theta}}s^{q-1}\leq C_{2}$
Now we return to (3.4) and we obtain
(3.15) $c_{k}^{\prime}\int_{\mathcal{C}_{\theta}}s^{1+\frac{k-1}{k}(q-1)}\phi^{\frac{k-1}{k}}\leq c_{k}\phi_{1}\int_{\mathcal{C}_{\theta}}s^{q-1}\leq C_{3}$
for some $C_{3}=C_{3}(n,k,p,\theta,\phi_{0},\phi_{1})$. Using $\phi\geq\phi_{0}$, this implies
(3.16) $\int_{\mathcal{C}_{\theta}}s^{\alpha(q)}\leq C_{4}$
for all $q\in[1,p]$, with $C_{4}$ depending only on $(n,k,p,\theta,\phi_{0},\phi_{1})$.
Since $\alpha(q)\geq 1$, (3.16) also yields a uniform $L^{1}$ bound for $s$:
(3.17) $\int_{\mathcal{C}_{\theta}}s\leq|\mathcal{C}_{\theta}|^{1-\frac{1}{\alpha(q)}}\Bigl(\int_{\mathcal{C}_{\theta}}s^{\alpha(q)}\Bigr)^{\frac{1}{\alpha(q)}}\leq C_{5}$
for all $q\in[1,p]$, with $C_{5}$ depending only on $(n,k,p,\theta,\phi_{0},\phi_{1})$.
Finally, the argument in the proof of [HIS25, Lem. 4.6] implies
(3.18) $s\leq C_{0}\quad\text{in }\mathcal{C}_{\theta}$
for all $q\in[1,p]$ with $C_{0}=C_{0}(n,k,p,\theta,\phi_{0},\phi_{1})$. This completes the proof. ∎
Proposition 3.2.
Let $\tilde{s}:=s/\ell$ where $s$ solves (3.1). Then
(3.19) $\sigma_{k}\Bigl(\ell\,\nabla^{2}\tilde{s}+\nabla\tilde{s}\otimes\nabla\ell+\nabla\ell\otimes\nabla\tilde{s}+\tilde{s}g,g\Bigr)=(\tilde{s}\ell)^{q-1}\phi\quad\text{in }\mathcal{C}_{\theta}$
and $\nabla_{\mu}\tilde{s}=0$ on $\partial\mathcal{C}_{\theta}$.
Lemma 3.3.
Let $s$ solve (3.1). Then
(3.20) $\max_{\mathcal{C}_{\theta}}s\geq\left(\frac{\phi_{0}}{\binom{n}{k}}\right)^{\frac{1}{k+1-p}}(1-\cos\theta)^{\frac{k}{k+1-p}}.$
Proof.
Let $\zeta_{\ast}\in\mathcal{C}_{\theta}$ be a maximum point of $\tilde{s}$. Then $\nabla\tilde{s}(\zeta_{\ast})=0$ and $\nabla^{2}\tilde{s}(\zeta_{\ast})\leq 0$. At $\zeta_{\ast}$,
(3.21) $\tau[s](\zeta_{\ast})=\ell(\zeta_{\ast})\nabla^{2}\tilde{s}(\zeta_{\ast})+\tilde{s}(\zeta_{\ast})g\leq\tilde{s}(\zeta_{\ast})g.$
Hence
(3.22) $\sigma_{k}(\tau[s](\zeta_{\ast}))\leq\sigma_{k}(\tilde{s}(\zeta_{\ast})g)=\binom{n}{k}\tilde{s}(\zeta_{\ast})^{k}.$
Using (3.1) and $\phi\geq\phi_{0}$, we obtain
(3.23) $\phi_{0}\leq(\tilde{s}(\zeta_{\ast})\ell(\zeta_{\ast}))^{1-q}\sigma_{k}(\tau[s](\zeta_{\ast}))\leq\binom{n}{k}\tilde{s}(\zeta_{\ast})^{k+1-q}\ell(\zeta_{\ast})^{1-q}.$
Thus, by $1-\cos\theta\leq\ell$ we get
(3.24) $s(\zeta_{\ast})\geq\left(\frac{\phi_{0}}{\binom{n}{k}}\right)^{\frac{1}{k+1-q}}(1-\cos\theta)^{\frac{k}{k+1-q}}.$
Finally, (3.20) follows from $0<\phi_{0}<1$, $0<1-\cos\theta<1$ and $q\in[1,p]$ with $1<p<k+1$. ∎
3.1. Rotationally symmetric hypersurfaces
Define
(3.25) $r_{\mathrm{out}}=\max_{x^{\prime}\in\Omega}|x^{\prime}|,\quad r_{\mathrm{in}}=\min_{x^{\prime}\in\partial\Omega}|x^{\prime}|.$
Assume $\det D^{2}f\geq\Lambda$ in $\Omega$ and $f=0$ on $\partial\Omega$. Consider the quadratic barrier
(3.26) $Q(x^{\prime})=\frac{\Lambda^{1/n}}{2}(|x^{\prime}|^{2}-r_{\mathrm{in}}^{2}),\quad x^{\prime}\in\Omega.$
Then $Q\geq f$ on $\partial\Omega$ and $\det D^{2}Q\leq\det D^{2}f$ in $\Omega$. By comparison principle,
(3.27) $Q\geq f\quad\text{in }\Omega\quad\implies\quad H\geq\frac{\Lambda^{1/n}}{2}r_{\mathrm{in}}^{2},$
where $H=-\min f=-f(0)$.
Recall that the Gauss curvature of $\Sigma$ is given by
(3.28) $\mathcal{K}=\frac{\det D^{2}f}{(1+|Df|^{2})^{(n+2)/2}}\quad\text{in }\Omega.$
Since $\phi=s^{1-q}\sigma_{k}\geq c_{k}s^{1-q}\sigma_{n}^{k/n}$ in $\mathcal{C}_{\theta}$ with $c_{k}=\binom{n}{k}$, we have
(3.29) $\mathcal{K}\geq c_{k}^{n/k}\phi^{-n/k}s^{n(1-q)/k}$
and
(3.30) $\det D^{2}f\geq c_{k}^{n/k}\phi_{1}^{-n/k}(s_{\max})^{\frac{n(1-q)}{k}},$
where $\phi_{0}\leq\phi\leq\phi_{1}$ with the constants $0<\phi_{0}<1<\phi_{1}$.
Theorem 3.4.
Let $\Sigma$ be a rotationally symmetric, strictly convex, $\theta$-capillary hypersurface whose capillary support function $s$ satisfies (3.1). Then
(3.31) $H\geq H_{\star},\quad H_{\star}=H_{\star}(n,k,p,\theta,\phi_{0},\phi_{1}).$
In particular, $H_{\star}\cos\theta\leq s\leq C_{0}$.
Proof.
The upper bound $s\leq C_{0}$ was established in 3.1. Due to (3.30) and (3.27), we have
(3.32) $H\geq\frac{c_{k}^{1/k}}{2}\phi_{1}^{-1/k}C_{0}^{\frac{1-q}{k}}r_{\mathrm{in}}^{2}\geq\frac{c_{k}^{1/k}}{2}\phi_{1}^{-1/k}C_{0}^{\frac{1-p}{k}}r_{\mathrm{in}}^{2},$
where we used that $C_{0}>1$ and $q\in[1,p]$. Since $\Sigma$ is rotationally symmetric, $r_{\mathrm{in}}=r_{\mathrm{out}}$ and thus $s\leq r_{\mathrm{in}}+H$. Now, by 3.3 and 2.8,
(3.33) $c_{0}\leq s_{\max}\leq r_{\mathrm{in}}+H\leq(1+\tan\theta)r_{\mathrm{in}},$
where $c_{0}=\left(\frac{\phi_{0}}{\binom{n}{k}}\right)^{\frac{1}{k+1-p}}(1-\cos\theta)^{\frac{k}{k+1-p}}$. Hence
(3.34) $r_{\mathrm{in}}\geq\frac{c_{0}}{1+\tan\theta},$
and the lower bound on $H$ follows. Due to $s\geq H\cos\theta$, the proof is complete. ∎
3.2. Even hypersurfaces
The argument in 3.4 uses the capillary $L_{p}$-Christoffel-Minkowski equation mainly through the inequality $\det D^{2}f\geq\Lambda$ for the Monge-Ampère measure of the graph function. Taken in isolation, this scalar inequality does not exclude degeneration of the base domain $\Omega$, and within this framework one cannot obtain a uniform positive lower bound for $H$ without an additional geometric input such as the rotationally symmetric assumption in conjunction with the capillarity assumption.
For general even, strictly convex, $\theta$-capillary hypersurfaces we keep the full equation and work directly at the level of area measures. From a sequence with $H_{i}\to 0$ we extract, by Blaschke’s selection theorem, a nontrivial limit body $K_{\infty}\subset e_{n+1}^{\perp}$ with linear span $L=\operatorname{lin}(K_{\infty})$, $\dim L=m\in\{1,\dots,n\}$. Using 3.5 and 3.11 we describe $S_{k}(K_{\infty},\cdot)$ on belts $\mathcal{B}\Subset\mathbb{S}^{n}_{\theta}$ at positive distance from $\mathbb{S}^{n}\cap L^{\perp}$. The measure identity together with $0<\phi_{0}\leq\phi\leq\phi_{1}$ yields a uniform positive lower bound for the $h_{i}^{1-p}$-weighted curvature on each such belt, whereas for a body contained in $L$ the structure of $S_{k}$ forces these contributions to vanish (or tend to zero) as the belt shrinks. This contradiction rules out $H_{i}\to 0$ and gives the desired uniform height lower bound in the general even case.
[…]
Theorem 3.5 ([GKW11], Thm. 6.2).
Let $L\subset\mathbb{R}^{n+1}$ be a linear subspace with $\dim L=m$ and $1\leq m\leq n$. Let $K\subset L$ be a convex body (with nonempty interior in $L$) and $k\in\{1,\dots,m-1\}$. Then, for every nonnegative measurable function $\psi$ on $\mathbb{S}^{n}$,
(3.35) $\int_{\mathbb{S}^{n}}\psi(u)\,dS_{k}(K,u)=c_{m,k}\int_{\mathbb{S}^{m-1}\cap L}I(\xi)\,dS_{k}^{L}(K,\xi),$
where
(3.36) $I(\xi)=\int_{\mathbb{S}^{n-m}\cap L^{\perp}}\int_{0}^{\pi/2}\psi(\sin\beta\,\xi+\cos\beta\,\eta)\sin^{m-k-1}\beta\,\cos^{n-m}\beta\,d\beta\,d\eta,$
and
(3.37) $c_{m,k}=\frac{\binom{m-1}{k}}{\binom{n}{k}}$.
Proof.
The integral formulation follows directly from [GKW11, Thm. 6.2], which states:
(3.38) $\binom{m-1}{k}\pi_{L,-k}^{*}S_{k}^{L}(K,\cdot)=\binom{n}{k}S_{k}(K,\cdot).$
By the definition of the lifting operator $\pi_{L,-k}^{*}$ (cf. [GKW11, Def. 5.2]):
(3.39) $\pi_{L,-k}^{*}S_{k}^{L}(K,A)=\int_{\mathbb{S}^{m-1}\cap L}\int_{H^{n+1-m}(L,\xi)\cap A}\langle\xi,w\rangle^{m-k-1}dw\,S_{k}^{L}(K,d\xi).$
In our coordinates, $w=\cos\beta\,\eta+\sin\beta\,\xi$, so $\langle\xi,w\rangle=\sin\beta$. Moreover, on the relatively open (n+1m)(n+1-m)-dimensional half-sphere
(3.40) $H^{n+1-m}(L,\xi)=\bigl\{w\in\mathbb{S}^{n}\setminus L^{\perp}:\operatorname{pr}_{L}(w)=\xi\bigr\}.$
we have $dw:=d\mathcal{H}^{n+1-m}(w)=\cos^{n-m}\beta\,d\beta\,d\mathcal{H}^{n-m}(\eta)$. Here, $\operatorname{pr}_{L}(w)$ is the spherical projection of $w$ on $\mathbb{S}^{n}\cap L$. ∎
Lemma 3.6.
Let $L\subset\mathbb{R}^{n+1}$ be a linear subspace with $\dim L=m\in\{1,\dots,n\}$, and let $K\subset L$ be a convex body (with nonempty interior) in $L$. Suppose $k\in\{1,\dots,m-1\}$. Let $\mathcal{U}\subset\mathbb{S}^{m-1}\cap L$ and $\mathcal{V}\subset\mathbb{S}^{n-m}\cap L^{\perp}$ be (relatively) open spherical caps with
(3.41) $S_{k}^{L}(K,\mathcal{U})>0\quad\text{and}\quad\mathcal{H}^{n-m}(\mathcal{V})>0.$
For angles $0<\beta_{1}<\beta_{2}<\pi/2$, define the belt
(3.42) $\mathcal{B}=\bigl\{u=\sin\beta\,\xi+\cos\beta\,\eta:\ \eta\in\mathcal{V},\ \xi\in\mathcal{U},\ \beta\in(\beta_{1},\beta_{2})\bigr\}\subset\mathbb{S}^{n}.$
Then
(3.43) $S_{k}(K,\mathcal{B})=c_{m,k}\mathcal{H}^{n-m}(\mathcal{V})S_{k}^{L}(K,\mathcal{U})\int_{\beta_{1}}^{\beta_{2}}\sin^{m-k-1}\beta\,\cos^{n-m}\beta\,d\beta.$
Proof.
The claim follows from 3.5 with the choice $\psi=\mathbf{1}_{\mathcal{B}}$. ∎
Lemma 3.7.
Let $K_{i}\subset\mathbb{R}^{n+1}$ be a sequence of origin-symmetric convex bodies with $K_{i}\to K_{\infty}$ in the Hausdorff metric and assume that $K_{\infty}\subset e_{n+1}^{\perp}$ is not a single point. Let
(3.44) $L=\operatorname{lin}(K_{\infty})\subset e_{n+1}^{\perp},\quad m=\dim L\in\{1,\dots,n\},\quad\mathcal{U}=\mathbb{S}^{m-1}\cap L.$
Then there exist constants $c_{\star}>0$ and $i_{0}\in\mathbb{N}$, angles $0<\beta_{1}<\beta_{2}<\theta$, and an open spherical cap $\mathcal{V}\subset\mathbb{S}^{n-m}\cap L^{\perp}$ centered at $e_{n+1}$, such that for the belt
(3.45) $\mathcal{B}=\bigl\{u=\sin\beta\,\xi+\cos\beta\,\eta:\xi\in\mathcal{U},\eta\in\mathcal{V},\beta\in(\beta_{1},\beta_{2})\bigr\}\subset\mathbb{S}^{n},$
the following hold:
  1. (i)
    $\mathcal{B}\Subset\operatorname{int}(\mathbb{S}^{n}_{\theta})$ and $\overline{\mathcal{B}}\cap(\mathbb{S}^{n}\cap L^{\perp})=\emptyset$;
  2. (ii)
    for all $i\geq i_{0}$ and all $u\in\overline{\mathcal{B}}$,
    (3.46) $h_{K_{i}}(u)\geq c_{\star}\,\sin\beta_{1}.$
Proof.
Write $h_{i}:=h_{K_{i}}$ and $h_{\infty}:=h_{K_{\infty}}$. Since $K_{\infty}$ has nonempty interior in $L$, there exists $c_{\star}>0$ such that
(3.47) $h_{\infty}(\xi)\geq 4c_{\star}\quad\forall\,\xi\in\mathcal{U}.$
By the uniform convergence of $h_{i}\to h_{\infty}$, there exists $i_{0}$ such that for all $i\geq i_{0}$,
(3.48) $h_{i}(\xi)\geq 2c_{\star}\quad\forall\,\xi\in\mathcal{U}.$
Since $K_{\infty}\subset L$, we have $h_{\infty}(\eta)=0$ for every $\eta\in\mathbb{S}^{n}\cap L^{\perp}$. Let $0<\beta_{1}<\beta_{2}<\theta$. Choose $\epsilon>0$ so small that $\epsilon<\theta-\beta_{2}$ and define the spherical cap $\mathcal{V}\subset\mathbb{S}^{n-m}\cap L^{\perp}$ by
(3.49) $\mathcal{V}=\{\eta\in\mathbb{S}^{n-m}\cap L^{\perp}:\ \angle(\eta,e_{n+1})<\epsilon\}.$
Then for any $\eta\in\overline{\mathcal{V}}$ and any $\beta\in[\beta_{1},\beta_{2}]$ we have
(3.50) $\langle\sin\beta\,\xi+\cos\beta\,\eta,\ e_{n+1}\rangle=\cos\beta\,\langle\eta,e_{n+1}\rangle\geq\cos\beta\,\cos\epsilon\geq\cos(\beta+\epsilon)>\cos\theta,$
so $\overline{\mathcal{B}}\subset\operatorname{int}(\mathbb{S}^{n}_{\theta})$. Also, since $\beta\geq\beta_{1}>0$, the set $\overline{\mathcal{B}}$ is disjoint from $\mathbb{S}^{n}\cap L^{\perp}$.
Next, since $h_{\infty}\equiv 0$ on $\mathbb{S}^{n}\cap L^{\perp}$, uniform convergence of $h_{i}\to h_{\infty}$ implies (after increasing $i_{0}$ if necessary) that for all $i\geq i_{0}$,
(3.51) $\sup_{\eta\in\overline{\mathcal{V}}}h_{i}(\eta)\leq c_{\star}\tan\beta_{1}.$
Let $i\geq i_{0}$ and $u\in\overline{\mathcal{B}}$. Then $u=\sin\beta\,\xi+\cos\beta\,\eta$ for some $\xi\in\overline{\mathcal{U}}$, $\eta\in\overline{\mathcal{V}}$, $\beta\in[\beta_{1},\beta_{2}]$. Choose $x_{i}\in K_{i}$ with $\langle x_{i},\xi\rangle=h_{i}(\xi)$. Since $K_{i}$ is origin-symmetric, we have $\langle x_{i},\eta\rangle\geq-h_{i}(\eta)$, hence
(3.52) $h_{i}(u)\geq\langle x_{i},u\rangle=\sin\beta\,h_{i}(\xi)+\cos\beta\,\langle x_{i},\eta\rangle\geq\sin\beta\,h_{i}(\xi)-\cos\beta\,h_{i}(\eta).$
Using (3.48), (3.51), and $\sin\beta\geq\sin\beta_{1}$, $\cos\beta\leq\cos\beta_{1}$, we obtain
(3.53) $h_{i}(u)\geq\sin\beta_{1}\,(2c_{\star})-\cos\beta_{1}\left(c_{\star}\tan\beta_{1}\right)=c_{\star}\sin\beta_{1},$
which proves (3.46). ∎
Theorem 3.8.
Suppose $\Sigma$ is an even, strictly convex, $\theta$-capillary hypersurface whose capillary support function $s$ satisfies (3.1). Then
(3.54) $H=\max_{x\in\Sigma}\langle x,e_{n+1}\rangle\geq H_{\star}>0,\quad H_{\star}\cos\theta\leq s\leq C_{0}$
with $H_{\star}=H_{\star}(k,p,\theta,\phi_{0},\phi_{1},C_{0})$.
Proof.
Let $K$ denote the union of $\widehat{\Sigma}$ and its reflection across the hyperplane $\{x_{n+1}=0\}$ and set $h:=h_{K}$. Assume for contradiction that there exist a sequence $(q_{i},\psi_{i},\Sigma_{i},s_{i},K_{i},h_{i})$ satisfying (3.1) with $\phi=\psi_{i}$, $q_{i}\in[1,p]$ and $\phi_{0}\leq\psi_{i}\leq\phi_{1}$, while
(3.55) $H_{i}=s_{i}((1-\cos\theta)e_{n+1})\to 0,\quad q_{i}\to q_{\ast}\in[1,p].$
Note that by 3.1, we have
(3.56) $\sup_{\mathcal{C}_{\theta}}s_{i}\leq C_{0}\quad\text{for all }i.$
In view of [HIS25, Lem. 4.2] and the Blaschke selection theorem, after passing to a subsequence, $K_{i}\to K_{\infty}$ in the Hausdorff metric. Then $K_{\infty}\subset e_{n+1}^{\perp}$ is origin-symmetric and it is not a point (by 3.3).
Let $L:=\operatorname{lin}(K_{\infty})$ and $m:=\dim L\in\{1,\dots,n\}$. Applying 3.7, we find $\mathcal{B}\Subset\operatorname{int}(\mathbb{S}^{n}_{\theta})$ and constants $c_{\star}>0$, $i_{0}\in\mathbb{N}$, and $0<\beta_{1}<\beta_{2}<\theta$ such that for all $i\geq i_{0}$ and all $u\in\overline{\mathcal{B}}$,
(3.57) $h_{i}(u)\geq c_{\star}\sin\beta_{1}.$
Since $\beta_{1}$ can be chosen so that $c_{\star}\sin\beta_{1}<1$, and $q_{i}\in[1,p]$, we obtain on $\overline{\mathcal{B}}$:
(3.58) $C_{0}^{1-p}\leq h_{i}^{1-q_{i}}(u)\leq(c_{\star}\sin\beta_{1})^{1-p}\quad\text{for all }u\in\overline{\mathcal{B}},\ i\geq i_{0}.$
Next, note that $m\geq k$. Otherwise, if $m<k$, by 3.9, then we have
(3.59) $S_{k}(K_{\infty},\overline{\mathcal{B}})=0.$
Since $S_{k}(K_{i},\cdot)\to S_{k}(K_{\infty},\cdot)$, it follows that
(3.60) $S_{k}(K_{i},\overline{\mathcal{B}})\to 0,$
and by (3.58),
(3.61) $\int_{\mathcal{B}}h_{i}^{1-q_{i}}\,dS_{k}(K_{i},u)\leq(c_{\star}\sin\beta_{1})^{1-p}S_{k}(K_{i},\mathcal{B})\to 0.$
On the other hand, we have
(3.62) $\binom{n}{k}\int_{\mathcal{B}}h_{i}^{1-q_{i}}\,dS_{k}(K_{i},u)=\int_{T^{-1}\mathcal{B}}\psi_{i}\geq\phi_{0}\,\mathcal{H}^{n}(\mathcal{B})>0,$
a contradiction. Thus $m\geq k$.
Case 1: $m\geq k+1$. Recall that $K_{\infty}$ has non-empty interior in $L$, so for $\mathcal{U}=\mathbb{S}^{m-1}\cap L$ we have $S^{L}_{k}(K_{\infty},\mathcal{U})>0$. Choose $\beta_{1},\beta_{2}$ as in 3.7. Then by 3.6,
(3.63) $S_{k}(K_{\infty},\mathcal{B})=c_{m,k}\mathcal{H}^{n-m}(\mathcal{V})S_{k}^{L}(K_{\infty},\mathcal{U})\int_{\beta_{1}}^{\beta_{2}}\sin^{m-k-1}\beta\,\cos^{n-m}\beta\,d\beta.$
Using (3.58), we obtain for $i\geq i_{0}$,
(3.64) $\int_{\mathcal{B}}h_{i}^{1-q_{i}}\,dS_{k}(K_{i},u)\geq C_{0}^{1-p}S_{k}(K_{i},\mathcal{B}).$
Taking $\liminf$ and using the weak convergence of $S_{k}(K_{i},\cdot)$,
(3.65) $\liminf_{i\to\infty}\int_{\mathcal{B}}h_{i}^{1-q_{i}}\,dS_{k}(K_{i},u)$
$\geq$ $C_{0}^{1-p}S_{k}(K_{\infty},\mathcal{B})$
$=$ $C_{0}^{1-p}c_{m,k}\mathcal{H}^{n-m}(\mathcal{V})S_{k}^{L}(K_{\infty},\mathcal{U})\int_{\beta_{1}}^{\beta_{2}}\sin^{m-k-1}\beta\,\cos^{n-m}\beta\,d\beta.$
On the other hand, we have
(3.66) $\binom{n}{k}\int_{\mathcal{B}}h_{i}^{1-q_{i}}\,dS_{k}(K_{i},u)$ $=\int_{T^{-1}\mathcal{B}}\psi_{i}$
$\leq\phi_{1}\,\mathcal{H}^{n}(\mathcal{B})$
$=\phi_{1}\,\mathcal{H}^{m-1}(\mathcal{U})\mathcal{H}^{n-m}(\mathcal{V})\int_{\beta_{1}}^{\beta_{2}}\sin^{m-1}\beta\,\cos^{n-m}\beta\,d\beta.$
Since the right-hand side is independent of $i$, we have
(3.67) $\binom{n}{k}\limsup_{i\to\infty}\int_{\mathcal{B}}h_{i}^{1-q_{i}}\,dS_{k}(K_{i},u)$
$\leq\phi_{1}\,\mathcal{H}^{m-1}(\mathcal{U})\mathcal{H}^{n-m}(\mathcal{V})\int_{\beta_{1}}^{\beta_{2}}\sin^{m-1}\beta\,\cos^{n-m}\beta\,d\beta.$
Combining the upper and lower bounds and cancelling the common factor $\mathcal{H}^{n-m}(\mathcal{V})$ we obtain
(3.68) $\frac{\binom{m-1}{k}C_{0}^{1-p}}{\phi_{1}}S_{k}^{L}(K_{\infty},\mathcal{U})\leq\mathcal{H}^{m-1}(\mathcal{U})\,\frac{\int_{\beta_{1}}^{\beta_{2}}\sin^{m-1}\beta\,\cos^{n-m}\beta\,d\beta}{\int_{\beta_{1}}^{\beta_{2}}\sin^{m-k-1}\beta\,\cos^{n-m}\beta\,d\beta}.$
Letting $\beta_{2}\downarrow\beta_{1}$ we get
(3.69) $\frac{\binom{m-1}{k}C_{0}^{1-p}}{\phi_{1}}S_{k}^{L}(K_{\infty},\mathcal{U})\leq\mathcal{H}^{m-1}(\mathcal{U})(\sin\beta_{1})^{k}.$
Letting $\beta_{1}\downarrow 0$ forces the right-hand side to tend to 0. This is a contradiction.
Case 2: $m=k$. In this case, by 3.11 (applied after approximating $K_{\infty}$ by polytopes) we have
(3.70) $S_{k}(K_{\infty},\omega)=0$
for every Borel set $\omega\subset\mathbb{S}^{n}$ with $\omega\cap(\mathbb{S}^{n}\cap L^{\perp})=\emptyset$. In particular, since $\overline{\mathcal{B}}\cap(\mathbb{S}^{n}\cap L^{\perp})=\emptyset$, we have
(3.71) $S_{k}(K_{\infty},\overline{\mathcal{B}})=0.$
Using (3.58) and weak convergence again, we get
(3.72) $S_{k}(K_{i},\mathcal{B})\to 0,$
and hence
(3.73) $\int_{\mathcal{B}}h_{i}^{1-q_{i}}\,dS_{k}(K_{i},u)\leq(\sup_{\mathcal{B}}h_{i}^{1-q_{i}})\,S_{k}(K_{i},\mathcal{B})\leq(c_{\star}\sin\beta_{1})^{1-p}S_{k}(K_{i},\mathcal{B})\to 0.$
On the other hand,
(3.74) $\binom{n}{k}\int_{\mathcal{B}}h_{i}^{1-q_{i}}\,dS_{k}(K_{i},u)=\int_{T^{-1}\mathcal{B}}\psi_{i}\geq\phi_{0}\,\mathcal{H}^{n}(\mathcal{B})>0,$
a contradiction.
Thus in all cases our assumption $H_{i}\to 0$ leads to a contradiction. Therefore there exists $H_{\star}>0$, depending only on $(n,k,p,\theta,\phi_{0},\phi_{1},C_{0})$, such that
(3.75) $H\geq H_{\star}$
for every even, strictly convex, $\theta$-capillary solution of (3.1) with $q\in[1,p]$ and $\phi_{0}\leq\phi\leq\phi_{1}$.
Finally, since $\Sigma$ is even, we have $H_{\star}e_{n+1}\in\widehat{\Sigma}$, and hence
(3.76) $s\geq H_{\star}\cos\theta.$
Remark 3.9.
Let $K\subset\mathbb{R}^{n+1}$ be a non-empty convex set and $L=\operatorname{lin}(K)$. Assume that $k>m=\dim L$. We show that $S_{k}(K,\cdot)\equiv 0$. For a Borel set $\omega\subset\mathbb{S}^{n}$ and $\rho>0$ define
(3.77) $B_{\rho}(K,\omega)=\{x\in\mathbb{R}^{n+1}:0<d(K,x)\leq\rho,\,u(K,x)\in\omega\},$
where $d(K,x)$ is the Euclidean distance from $x$ to $K$, $p(K,x)$ is a nearest point of $K$ to $x$, and
(3.78) $u(K,x)=\frac{x-p(K,x)}{\lvert x-p(K,x)\rvert}$.
By the local Steiner formula (cf. [Sch14, (4.13)]),
(3.79) $\mathcal{H}^{n+1}\bigl(B_{\rho}(K,\omega)\bigr)=\frac{1}{n+1}\sum_{j=0}^{n}\binom{n+1}{j}\rho^{n+1-j}S_{j}(K,\omega).$
Since $K\subset L$, we have
(3.80) $\{x\in\mathbb{R}^{n+1}:d(K,x)\leq\rho\}\subset(K+\rho B_{L})+\rho B_{L^{\perp}},$
where $B_{L}=\mathbb{B}\cap L$ and $B_{L^{\perp}}=\mathbb{B}\cap L^{\perp}$ are the unit balls in $L$ and $L^{\perp}$, respectively. In particular, for $\rho\leq 1$:
(3.81) $\mathcal{H}^{n+1}\bigl(B_{\rho}(K,\omega)\bigr)\leq\mathcal{H}^{m}(K+\rho B_{L})\,\mathcal{H}^{n+1-m}(\rho B_{L^{\perp}})\leq C\rho^{n+1-m},$
where $C:=\mathcal{H}^{m}(K+B_{L})\,\mathcal{H}^{n+1-m}(B_{L^{\perp}})$.
On the other hand, if $S_{k}(K,\omega)>0$ for some Borel set $\omega$, then (3.79) yields
(3.82) $\mathcal{H}^{n+1}\bigl(B_{\rho}(K,\omega)\bigr)\geq c\rho^{n+1-k},\quad c=c(n,K,\omega).$
Combining (3.81) and (3.82) gives
(3.83) $c\rho^{n+1-k}\leq C\,\rho^{n+1-m}\qquad\text{for all }0<\rho\leq 1.$
Since $k>m$, we get a contradiction by letting $\rho\to 0$.
Lemma 3.10 ([Sch14], p. 216).
Let $d\geq 2$ and let $P\subset\mathbb{R}^{d}$ be a (not necessarily full-dimensional) convex polytope. For $k\in\{0,1,\dots,d-1\}$ and every Borel set $\omega\subset\mathbb{S}^{d-1}$,
(3.84) $S_{k}(P,\omega)=\sum_{F\in\mathcal{F}_{k}(P)}\frac{\mathcal{H}^{d-1-k}\big(N(P,F)\cap\omega\big)}{\omega_{d-k}}\;\mathcal{H}^{k}(F).$
Here $\mathcal{F}_{k}(P)$ is the set of $k$-faces of $P$, $N(P,F)$ is the normal cone of $P$ at $F$ (i.e. the set of all outer normal vectors of $K$ at any $x\in\operatorname{relint}F$ together with the zero vector), $\omega_{m}=\mathcal{H}^{m}(\mathbb{S}^{m})$, and $S_{k}(P,\cdot)$ is the $k$-th area measure of $P$ on $\mathbb{S}^{d-1}$. In particular,
(3.85) $\operatorname{supp}S_{k}(P,\cdot)\ \subset\ \bigcup_{F\in\mathcal{F}_{k}(P)}\big(N(P,F)\cap\mathbb{S}^{d-1}\big)\ =\ \bigcup_{F\in\mathcal{F}_{k}(P)}\nu_{P}(\operatorname{relint}(F)),$
where $\nu_{P}$ denotes the spherical image of $P$.
Corollary 3.11.
Suppose $L\subset\mathbb{R}^{d}$ is a linear subspace with $m=\dim L\in\{1,\dots,d-1\}$, and let $P\subset L$ be an $m$-dimensional polytope. Then $S_{m}(P,\cdot)$ is concentrated on $\mathbb{S}^{d-1}\cap L^{\perp}$:
(3.86) $S_{m}(P,\omega)$ $=\frac{\mathcal{H}^{d-1-m}\big(L^{\perp}\cap\omega\big)}{\omega_{d-m}}\;\mathcal{H}^{m}(P),\quad\operatorname{supp}S_{m}(P,\cdot)$ $\subset\ \mathbb{S}^{d-1}\cap L^{\perp}.$
Proof.
For $k=m$, the only $m$-face is $P$ and $N(P,P)=L^{\perp}$. ∎
4. Regularity Estimates
Lemma 4.1.
Suppose $\Sigma$ is an even, strictly convex, $\theta$-capillary hypersurface whose capillary support function $s$ satisfies (3.1). Then
(4.1) $\sigma_{1}(\tau^{\sharp}[s])\leq C\quad\text{in }\mathcal{C}_{\theta},$
for some constant $C$ depending only on $n,k,p,\theta,\phi$.
Proof.
Let $F=\sigma_{k}^{\frac{1}{k}}$. Then
(4.2) $F(\tau^{\sharp}[s])=s^{\frac{q-1}{k}}\phi^{\frac{1}{k}}.$
Using the identity
(4.3) $\nabla^{2}_{ii}\sigma_{1}=\Delta\tau_{ii}-n\tau_{ii}+\sigma_{1}$
and the concavity of $F$, there holds
(4.4) $F^{ij}g_{ij}\sigma_{1}\leq F^{ij}\nabla^{2}_{ij}\sigma_{1}+ns^{\frac{q-1}{k}}\phi^{\frac{1}{k}}-\Delta(s^{\frac{q-1}{k}}\phi^{\frac{1}{k}}).$
We calculate
(4.5) $-k\Delta(s^{\frac{q-1}{k}}\phi^{\frac{1}{k}})=$ $(1-q)s^{\frac{q-1}{k}-1}\phi^{\frac{1}{k}}\sigma_{1}-n(1-q)s^{\frac{q-1}{k}}\phi^{\frac{1}{k}}$
$+\frac{1}{k}(q-1)(k+1-q)s^{\frac{q-1}{k}-2}|\nabla s|^{2}\phi^{\frac{1}{k}}$
$+2(1-q)s^{\frac{q-1}{k}-1}\langle\nabla s,\nabla\phi^{\frac{1}{k}}\rangle-ks^{\frac{q-1}{k}}\Delta\phi^{\frac{1}{k}}.$
Due to the concavity of $F$, we have $\operatorname{tr}(\dot{F})\geq c_{k}$. By 3.8, we have
(4.6) $1/C_{1}\leq s\leq C_{1}$
for some constant $C_{1}>1$ depending only $n,k,p,\theta,\phi$. It follows from [HIS25, Lem. 4.8] that
(4.7) $|\nabla s|\leq\frac{C_{1}}{\sin\theta}.$
Hence, if $\sigma_{1}$ attains its maximum in the interior of $\mathcal{C}_{\theta}$, we have
(4.8) $c_{k}\sigma_{1}\leq$ $~nC_{1}\|\phi^{\frac{1}{k}}\|_{C^{0}}+nC_{1}\|\phi^{\frac{1}{k}}\|_{C^{0}}+C_{1}^{2}\big(\frac{C_{1}}{\sin\theta}\big)^{2}\|\phi^{\frac{1}{k}}\|_{C^{0}}$
$~+2\frac{C_{1}^{2}}{\sin\theta}\|\phi^{\frac{1}{k}}\|_{C^{1}}+C_{1}\|\phi^{\frac{1}{k}}\|_{C^{2}},$
where we also used that $q\in[1,p]$ with $1<p<k+1$. Thus we have
(4.9) $\sigma_{1}\leq C$
for some constant $C=C(n,k,p,\theta,\phi)$.
Now we need to treat the case that the maximum of $\sigma_{1}$ is attained at a boundary point, say $p_{\ast}$. Let $\{\mu\}\cup\{e_{\alpha}\}_{\alpha\geq 2}$ be an orthonormal basis of eigenvectors of $\tau^{\sharp}[s]$ at $p_{\ast}$ such that $\tau_{ij}=\lambda_{i}\delta_{ij}$. Moreover, using
(4.10) $\nabla_{\mu}\tau_{\alpha\beta}=\bigl(\tau_{\mu\mu}g_{\alpha\beta}-\tau_{\alpha\beta}\bigr)\cot\theta,\quad 2\leq\alpha,\beta\leq n,$
and $\nabla_{\mu}s=\cot\theta\,s$, we obtain at $p_{\ast}$ that
(4.11) $0\leq F^{\mu\mu}\nabla_{\mu}\sigma_{1}$ $\leq\cot\theta\Bigl((n+1)\phi^{\frac{1}{k}}s^{\frac{q-1}{k}}-F^{\mu\mu}\sigma_{1}-\sum_{i}F^{ii}\lambda_{\mu}\Bigr)$
$\quad+s^{\frac{q-1}{k}}\Bigl(\nabla_{\mu}\phi^{\frac{1}{k}}+\frac{q-1}{k}\cot\theta\,\phi^{\frac{1}{k}}\Bigr)$
and
(4.12) $\sigma_{1}$ $\leq\frac{s^{\frac{q-1}{k}}\max_{\mathcal{C}_{\theta}}\bigl|\nabla_{\mu}\phi^{\frac{1}{k}}\bigr|}{\cot\theta\,F^{\mu\mu}}+\Bigl(n+1+\frac{q-1}{k}\Bigr)\frac{s^{\frac{q-1}{k}}\phi^{\frac{1}{k}}}{F^{\mu\mu}}$
$\leq\frac{C_{1}\|\phi^{\frac{1}{k}}\|_{C^{1}}}{\cot\theta\,F^{\mu\mu}}+(n+2)\,\frac{C_{1}\|\phi^{\frac{1}{k}}\|_{C^{0}}}{F^{\mu\mu}},$
see [HIS25, (4.4),(4.5)] for details.
Next we show that $F^{\mu\mu}$ cannot be very small. By (4.6), we get
(4.13) $c_{1}=(\min_{\mathcal{C}_{\theta}}\phi)C_{1}^{1-p}$ $\leq\phi s^{q-1}=\sigma_{k}(\lambda)$
$=\lambda_{\mu}\sigma_{k-1}(\lambda|\lambda_{\mu})+\sigma_{k}(\lambda|\lambda_{\mu})$
$\leq\lambda_{\mu}\sigma_{k-1}(\lambda|\lambda_{\mu})+c_{1}\sigma_{k-1}(\lambda|\lambda_{\mu})^{\frac{k}{k-1}}$
$\leq c_{2}\lambda_{\mu}F^{\mu\mu}+c_{3}(F^{\mu\mu})^{\frac{k}{k-1}},$
where we used that
(4.14) $F^{\mu\mu}$ $=\frac{1}{k}\sigma_{k}^{\frac{1-k}{k}}(\lambda)\sigma_{k-1}(\lambda|\lambda_{\mu})$
$=\frac{1}{k}(s^{q-1}\phi)^{\frac{1-k}{k}}\sigma_{k-1}(\lambda|\lambda_{\mu})$
$\geq\frac{1}{k}C_{1}^{\frac{(p-1)(1-k)}{k}}\|\phi^{\frac{1}{k}}\|_{C^{0}}^{1-k}\sigma_{k-1}(\lambda|\lambda_{\mu}).$
Note that all these constants $c_{i}$ depend only on $n,p,k,\theta,\phi$.
Substituting (4.13) in (4.11), we obtain
(4.15) $0\leq F^{\mu\mu}\nabla_{\mu}\sigma_{1}\leq$ $~\Bigl((n+1)\phi^{\frac{1}{k}}s^{\frac{q-1}{k}}-\sum_{i}F^{ii}\lambda_{\mu}\Bigr)\cot\theta$
$~+s^{\frac{q-1}{k}}\Bigl(\nabla_{\mu}\phi^{\frac{1}{k}}+\frac{q-1}{k}\cot\theta\,\phi^{\frac{1}{k}}\Bigr)$
$\leq$ $~\frac{\cot\theta}{c_{2}}\sum_{i}F^{ii}\Bigl(-\frac{c_{1}}{F^{\mu\mu}}+c_{3}(F^{\mu\mu})^{\frac{1}{k-1}}\Bigr)$
$~+C_{1}\Bigl(\|\phi^{\frac{1}{k}}\|_{C^{1}}+\big(n+2\big)\cot\theta\,\|\phi^{\frac{1}{k}}\|_{C^{0}}\Bigr).$
Hence $F^{\mu\mu}$ cannot be small, and in view of (4.12), $\sigma_{1}$ is bounded above and the bound depends only on $n,p,k,\theta,\phi$. ∎
In view of 4.1, the higher-order regularity follows form [LT86] and Schauder estimate.
Proposition 4.2.
Suppose $\Sigma$ is an even, strictly convex, $\theta$-capillary hypersurface whose capillary support function $s$ satisfies (3.1). Then for any $m\geq 1$ we have $\|s\|_{C^{m}}\leq C_{m}$ for some constant depending only on $n,p,k,\theta,\phi$.
5. Strict Convexity
Theorem 5.1.
Let $\theta\in(0,\pi/2)$, $1\leq k<n$ and $q\geq 1$. Suppose $\phi\in C^{2}(\mathcal{C}_{\theta})$ satisfies
(5.1) $\nabla^{2}\phi^{-\frac{1}{q+k-1}}+g\,\phi^{-\frac{1}{q+k-1}}\;\geq 0\quad\text{in }\mathcal{C}_{\theta},$
and the boundary condition
(5.2) $\nabla_{\mu}\phi^{-\frac{1}{q+k-1}}\;\leq\;\cot\theta\,\phi^{-\frac{1}{q+k-1}}\quad\text{on }\partial\mathcal{C}_{\theta}.$
Let $0\leq s\in C^{2}(\mathcal{C}_{\theta})$ be a capillary function, i.e.
(5.3) $\nabla_{\mu}s=\cot\theta\,s\quad\text{on }\partial\mathcal{C}_{\theta},$
with
(5.4) $\tau^{\sharp}[s]\geq 0\quad\text{in }\mathcal{C}_{\theta},$
and suppose that $s$ solves
(5.5) $\sigma_{k}\bigl(\tau^{\sharp}[s]\bigr)=s^{q-1}\,\phi\quad\text{in }\mathcal{C}_{\theta}.$
Denote by $\lambda_{1}$ the smallest eigenvalue of $\tau^{\sharp}[s]$. If $s>0$, then $\lambda_{1}>0.$
Proof.
The argument is the same as in [HIS25, Thm. 3.1] for $q=1$. Define
(5.6) $F=\sigma_{k}^{1/k},\quad f=\bigl(s^{q-1}\phi\bigr)^{1/k}.$
When $\phi$ satisfies (5.1), we have in the interior of $\mathcal{C}_{\theta}$ that
(5.7) $L[\lambda_{1}]=F^{ij}\nabla^{2}_{ij}\lambda_{1}-c\bigl(\lambda_{1}+\lvert\nabla\lambda_{1}\rvert\bigr)\leq 0$
in the viscosity sense; for details see [BIS23a, Thm. 2.2] or [CH25, (3.20)]. Therefore, it suffices to carry out the boundary analysis in Step 1 of the proof of [HIS25, Thm. 3.1] at a point $p_{\ast}\in\partial\mathcal{C}_{\theta}$ where $\lambda_{1}(p_{\ast})=0$ while $\lambda_{1}>0$ in the interior of $\mathcal{C}_{\theta}$: we need a boundary condition on $\phi$ which ensures that
(5.8) $\tau_{ii}(p_{\ast})=0\implies\nabla_{\mu}\tau_{ii}(p_{\ast})\geq 0.$
Choose an orthonormal frame $\{e_{i}\}_{i=1}^{n}$ at $p_{\ast}$ such that
(5.9) $e_{1}=\mu,\quad e_{\alpha}\in T_{p_{\ast}}\partial\mathcal{C}_{\theta}\quad\text{ for }\alpha=2,\dots,n,$
and $\tau^{\sharp}[s]$ is diagonal in this frame at $p_{\ast}$ such that $\tau_{ij}=\lambda_{i}\,\delta_{ij}$.
For $i=\alpha\geq 2$, (5.8) follows directly from the boundary identity (4.10). For $i=1$, note that (5.5) is equivalent to
(5.10) $F\bigl(\tau^{\sharp}[s]\bigr)=f\quad\text{in }\mathcal{C}_{\theta}.$
Differentiating (5.10) in the $\mu$-direction gives
(5.11) $\sum_{i}F^{ii}\,\nabla_{\mu}\tau_{ii}=\nabla_{\mu}f.$
Using (4.10) for $\alpha\geq 2$, we obtain
(5.12) $F^{\mu\mu}\nabla_{\mu}\tau_{\mu\mu}=\nabla_{\mu}f+\sum_{\alpha\geq 2}F^{\alpha\alpha}(\tau_{\alpha\alpha}-\tau_{\mu\mu})\cot\theta.$
At $p_{\ast}$ we have $\tau_{\mu\mu}(p_{\ast})=\lambda_{1}(p_{\ast})=0$. By the 11-homogeneity of $F$,
(5.13) $\sum_{i}F^{ii}\tau_{ii}=F(\tau)=f,$
hence at $p_{\ast}$,
(5.14) $\sum_{\alpha\geq 2}F^{\alpha\alpha}\tau_{\alpha\alpha}=f.$
Evaluating (5.12) at $p_{\ast}$ yields
(5.15) $\nabla_{\mu}\tau_{\mu\mu}=\frac{\nabla_{\mu}f+f\cot\theta}{F^{\mu\mu}}.$
Thus (5.8) for $i=1$ holds provided that
(5.16) $\nabla_{\mu}\log f\geq-\cot\theta\quad\text{on }\partial\mathcal{C}_{\theta}.$
It remains to express (5.16) in terms of $\phi$. Since
(5.17) $f=s^{\frac{q-1}{k}}\phi^{\frac{1}{k}},\quad\text{and}\quad\nabla_{\mu}\log s=\cot\theta\quad\text{on }\partial\mathcal{C}_{\theta},$
we require that
(5.18) $\nabla_{\mu}\log\phi\geq-(k+q-1)\cot\theta\quad\text{on }\partial\mathcal{C}_{\theta},$
which is precisely (5.2). ∎
6. Existence and uniqueness
For $q\in[1,p]$ set
(6.1) $\phi_{q}=\phi^{\frac{q+k-1}{p+k-1}}.$
For $(q,s)$ with $q\in[1,p]$ and $s\in C^{l+2,\alpha}_{\mathrm{even}}(\mathcal{C}_{\theta})$ such that $s>0$ in $\mathcal{C}_{\theta}$ we define
(6.2) $\left\{\begin{aligned} F(q,s)&=\sigma_{k}(\tau^{\sharp}[s])-s^{q-1}\phi_{q}&&\text{in }\mathcal{C}_{\theta},\\ G(q,s)&=\nabla_{\mu}s-\cot\theta\,s&&\text{on }\partial\mathcal{C}_{\theta}.\end{aligned}\right$.
If $(F(q,s),G(q,s))=(0,0)$, then $s$ solves
(6.3) $\left\{\begin{aligned} \sigma_{k}(\tau^{\sharp}[s])&=s^{q-1}\phi_{q}&&\text{in }\mathcal{C}_{\theta},\\ \nabla_{\mu}s&=\cot\theta\,s&&\text{on }\partial\mathcal{C}_{\theta}.\end{aligned}\right$.
For $q=1$, by [HIS25, Thm. 1.2] there exists a unique even, smooth, strictly convex, $\theta$-capillary solution $s_{1}$ of this problem.
Assume now that $\phi_{0}\leq\phi\leq\phi_{1}$ with $0<\phi_{0}<1<\phi_{1}$. Then, for every $q\in[1,p]$,
(6.4) $\phi_{0}^{\frac{q+k-1}{p+k-1}}\;\leq\;\phi_{q}\;\leq\;\phi_{1}^{\frac{q+k-1}{p+k-1}}.$
Applying 3.1, 3.3 and 3.8 with $\phi$ replaced by $\phi_{q}$ we obtain constants $C_{0},c_{0}>0$, independent of $q$, such that every even solution $s$ of (6.3) with $\tau^{\sharp}[s]>0$ satisfies
(6.5) $c_{0}\leq s\leq C_{0}\quad\text{in }\mathcal{C}_{\theta}.$
Moreover, since $\phi_{q}$ satisfies the structural assumptions of 5.1, when $s$ is a solution of (6.3) with $\tau^{\sharp}[s]\geq 0$ and $s>0$, we must have
(6.6) $\tau^{\sharp}[s]>0\quad\text{in }\mathcal{C}_{\theta}.$
Combining (6.5) and (6.6) with 4.1 and 4.2, we obtain a uniform $C^{4,\alpha}$ bound: there exists $C>0$ such that
(6.7) $\|s\|_{C^{4,\alpha}(\mathcal{C}_{\theta})}\leq C$
for all even capillary solutions $s$ of (6.3) with $\tau^{\sharp}[s]>0$, uniformly in $q\in[1,p]$.
Let $R>C$ and define the bounded open set
(6.8) $\mathcal{O}=\bigl\{s\in C^{l+2,\alpha}_{\mathrm{even}}(\mathcal{C}_{\theta}):\|s\|_{C^{4,\alpha}(\mathcal{C}_{\theta})}<R,2s>c_{0},\ \tau^{\sharp}[s]>0\bigr\}.$
By (6.5), (6.6), and (6.7),
(6.9) $(F(q,s),G(q,s))\neq(0,0)\quad\text{for all }(q,s)\in[1,p]\times\partial\mathcal{O}.$
Therefore, by [LLN17, Thm. 1], for each $q\in[1,p]$ there is a well-defined integer-valued degree
(6.10) $d(q)=\deg\bigl((F(q,\cdot),G(q,\cdot)),\mathcal{O},0\bigr),$
which is homotopy invariant in $q$, in particular, $d(1)=d(p)$.
Due to [HIS25, Lem. 5.4], the linearized operator $\mathcal{L}:=D_{s}(F,G)(1,s_{1})$ has trivial kernel (the only kernel directions in the full class correspond to horizontal translations, which are odd). Together with 6.1 this implies that $\mathcal{L}$ is an isomorphism. Hence, from [LLN17, Thm. 1.1, Cor. 2.1] it follows that $d(1)=\pm 1$ and there exists $s\in\mathcal{O}$ such that
(6.11) $(F(p,s),G(p,s))=(0,0).$
Next, we establish uniqueness of solutions to (1.7) in the class of even, strictly convex capillary hypersurfaces.
Assume that $s_{1}$ and $s_{2}$ are two even, strictly convex capillary solutions to
(6.12) $\left\{\begin{aligned} \sigma_{k}(\tau^{\sharp}[s])&=s^{p-1}\phi&&\text{in }\mathcal{C}_{\theta},\\ \nabla_{\mu}s&=\cot\theta\,s&&\text{on }\partial\mathcal{C}_{\theta}.\end{aligned}\right$.
Using the mixed-volume interpretation of $\int_{\mathcal{C}_{\theta}}s\sigma_{k}(\tau^{\sharp}[\cdot])$ we obtain
(6.13) $\int_{\mathcal{C}_{\theta}}s_{2}s_{1}^{p-1}\phi$ $=\int_{\mathcal{C}_{\theta}}s_{2}\,\sigma_{k}(\tau^{\sharp}[s_{1}])$
$=(n+1)\binom{n}{k}V\bigl(s_{2},\underbrace{s_{1},\ldots,s_{1}}_{k\text{ copies}},\ell,\ldots,\ell\bigr)$
$\geq(n+1)\binom{n}{k}V\bigl(\underbrace{s_{1},\ldots,s_{1}}_{k+1\text{ copies}},\ell,\ldots,\ell\bigr)^{\frac{k}{k+1}}V\bigl(\underbrace{s_{2},\ldots,s_{2}}_{k+1\text{ copies}},\ell,\ldots,\ell\bigr)^{\frac{1}{k+1}}$
$=\Bigl(\int_{\mathcal{C}_{\theta}}s_{1}^{p}\phi\Bigr)^{\frac{k}{k+1}}\Bigl(\int_{\mathcal{C}_{\theta}}s_{2}^{p}\phi\Bigr)^{\frac{1}{k+1}},$
where we used Alexandrov-Fenchel’s inequality (see [MWWX25, Thm. 3.1]):
(6.14) $V(s_{1},s_{2},s_{3},\ldots,s_{n+1})^{2}\geq V(s_{1},s_{1},s_{3},\ldots,s_{n+1})\,V(s_{2},s_{2},s_{3},\ldots,s_{n+1}).$
On the other hand, by the Hölder inequality we have
(6.15) $\int_{\mathcal{C}_{\theta}}s_{2}s_{1}^{p-1}\phi\leq\Bigl(\int_{\mathcal{C}_{\theta}}s_{2}^{p}\phi\Bigr)^{\frac{1}{p}}\Bigl(\int_{\mathcal{C}_{\theta}}s_{1}^{p}\phi\Bigr)^{\frac{p-1}{p}}.$
Combining (6.13) and (6.15) yields
(6.16) $\Bigl(\int_{\mathcal{C}_{\theta}}s_{1}^{p}\phi\Bigr)^{\frac{k}{k+1}}\Bigl(\int_{\mathcal{C}_{\theta}}s_{2}^{p}\phi\Bigr)^{\frac{1}{k+1}}\leq\Bigl(\int_{\mathcal{C}_{\theta}}s_{2}^{p}\phi\Bigr)^{\frac{1}{p}}\Bigl(\int_{\mathcal{C}_{\theta}}s_{1}^{p}\phi\Bigr)^{\frac{p-1}{p}}.$
Rearranging, we obtain
(6.17) $\Bigl(\int_{\mathcal{C}_{\theta}}s_{1}^{p}\phi\Bigr)^{\frac{p-k-1}{p(k+1)}}\geq\Bigl(\int_{\mathcal{C}_{\theta}}s_{2}^{p}\phi\Bigr)^{\frac{p-k-1}{p(k+1)}}.$
Since $1<p<k+1$, we obtain
(6.18) $\int_{\mathcal{C}_{\theta}}s_{1}^{p}\phi\leq\int_{\mathcal{C}_{\theta}}s_{2}^{p}\phi.$
Interchanging $s_{1}$ and $s_{2}$ gives
(6.19) $\int_{\mathcal{C}_{\theta}}s_{2}^{p}\phi\leq\int_{\mathcal{C}_{\theta}}s_{1}^{p}\phi,$
so in fact
(6.20) $\int_{\mathcal{C}_{\theta}}s_{1}^{p}\phi=\int_{\mathcal{C}_{\theta}}s_{2}^{p}\phi.$
Thus equality holds in (6.13), and by the equality case in the Alexandrov-Fenchel inequality (6.14) we obtain $s_{1}=s_{2}$, since the equation is not scale invariant. This proves the uniqueness.
Lemma 6.1.
(6.21) $\mathcal{L}:C^{2,\alpha}_{\mathrm{even}}(\mathcal{C}_{\theta})\to C^{\alpha}_{\mathrm{even}}(\mathcal{C}_{\theta})\times C^{1,\alpha}_{\mathrm{even}}(\partial\mathcal{C}_{\theta})$
is an isomorphism.
Proof.
Since $\tau^{\sharp}[s_{1}]>0$, the matrix $a^{ij}]$ defined by $a^{ij}=\sigma_{k}^{ij}(\tau^{\sharp}[s_{1}])$ is uniformly positive definite on $\mathcal{C}_{\theta}$, and
(6.22) $Lv=a^{ij}\nabla^{2}_{ij}v+\mathrm{tr}(a)v,\quad v\in C^{2,\alpha}(\mathcal{C}_{\theta})$
is uniformly elliptic with $C^{\alpha}$ coefficients. Define the boundary operator
(6.23) $Mv=\nabla_{(-\mu)}v+\cot\theta\,v\quad\text{on }\partial\mathcal{C}_{\theta}.$
Using the stereographic projection from south pole $\Pi:\mathbb{S}^{n}\setminus\{-e_{n+1}\}\to\mathbb{R}^{n}$, $\Pi(x^{\prime},x_{n+1})=\frac{x^{\prime}}{1+x_{n+1}}$, we can rewrite $\mathcal{L}$ and $\mathcal{M}$ on $\Omega:=B_{\tan(\theta/2)}(0)\subset\mathbb{R}^{n}$:
[…]
with $a^{ij},b^{i},c\in C^{\infty}(\overline{\Omega}),\beta^{i}\in C^{\infty}(\partial\Omega)$, and $M$ uniformly oblique:
(6.25) $\langle\beta(x),-\frac{x}{|x|}\rangle=\frac{1}{1+\cos\theta}\quad\text{for all }x\in\partial\Omega.$
Therefore, by [Lie13, Thm. 2.30], the map
(6.26) $\mathcal{L}:C^{2,\alpha}(\mathcal{C}_{\theta})$ $\to C^{\alpha}(\mathcal{C}_{\theta})\times C^{1,\alpha}(\partial\mathcal{C}_{\theta}),$
$\mathcal{L}(v)$ $=\bigl(Lv,Mv\bigr),$
is a Fredholm operator of index 0. In particular, we have
(6.27) $\dim\ker\mathcal{L}=\dim\mathrm{coker}\,\mathcal{L}.$
Note that $\mathcal{L}$ preserves evenness and also under the stereographic projection from the south pole, evenness is preserved: if $v\circ\mathcal{R}=v$ on $\mathcal{C}_{\theta}$ and $u(x)=v(\Pi^{-1}(x))$, then u(x)=u(x)u(-x)=u(x) on $B_{\tan(\theta/2)}(0)$. By [HIS25, Lem. 5.4], if $v\in C_{\text{even}}^{2}(\mathcal{C}_{\theta})$ and satisfies
(6.28) $Lv=0\quad\text{in }\mathcal{C}_{\theta},\quad Mv=0\quad\text{on }\partial\mathcal{C}_{\theta},$
then $v\equiv 0$. In other words,
(6.29) $\ker\mathcal{L}\cap C^{2,\alpha}_{\mathrm{even}}(\mathcal{C}_{\theta})=\{0\}.$
Thus, in the even class, $\dim\mathrm{coker}\,\mathcal{L}=\dim\ker\mathcal{L}=0$ and $\mathcal{L}$ is an isomorphism. ∎
Remark 6.2.
We could also use the continuity method to solve the capillary even $L_{p}$-Christoffel–Minkowski problem. We may interpolate between $1$ and $\phi$ via the path
$H:[0,1]\to C^{\infty}(\mathcal{C}_{\theta}),\quad t\mapsto H(t,\cdot),$
defined by
[…]
where
$q(t):=1+(p-1)(2t-1),\quad t\in\bigl[\tfrac{1}{2},1\bigr].$
Then
(6.31) $H(0,\zeta)=1,\quad H\!\left(\tfrac{1}{2},\zeta\right)=\phi(\zeta)^{\frac{k}{p+k-1}},\quad H(1,\zeta)=\phi(\zeta).$
Now consider the equation
[…]
$\nabla_{\mu}s$ $=\cot\theta\,s.$
For $t=0$, the model capillary support function $\ell$ is the unique even solution of this problem. For every $t\in[0,1]$, the structural assumptions required by the constant rank theorem are satisfied. The closedness in the continuity method follows from the a priori estimates established above, while openness is as in the standard (closed) case.
Acknowledgments
Hu was supported by the National Key Research and Development Program of China (Grant No. 2021YFA1001800). Ivaki was supported by the Austrian Science Fund (FWF) under Project P36545.
We would like to thank Georg Hofstätter for helpful discussions.